### v1

# The Battle of the Water Futures

The Battle of the Water Futures
Dennis Zanutto    Christos Michalopoulos    Lydia Tsiami    Andre Artelt    Jasmin Brandt    Demetrios Eliades    Stelios Vrachimis    Stefano Alvisi    Valentina Marsili    Filippo Mazzoni    Panagiotis Samartzis    Barbara Hammer    Phoebe Koundouri    Marios Polycarpou    Dragan Savić
(November 2025)
1 Introduction
1.1 Disclaimer
The networks and problem presented in this competition are synthetic and do not represent any real-world situation. While we utilize real data from trusted sources where possible (e.g., CBS for municipal data, KNMI for wheater observations), other aspects rely on assumptions based on expert knowledge.
We are committed to the model’s integrity and, based on competitor feedback, we will continuously test and refine the model (e.g., balancing electricity and infrastructure costs) to ensure it remains physically sound and balanced. If you identify any significant inconsistencies, please inform us and we will review them promptly.
All versions are archived on Zenodo, with the latest version establishing the official standard. An open-source evaluator will also be released in the WaterBenchmarkHub (Artelt, 2025), allowing competitors to test their solutions before the final submission.
1.2 Background and motivation
Water distribution systems (WDS) grow organically with cities and are therefore inherently subject to deep uncertainties. While secondary and tertiary network components can still be adequately selected using established engineering principles, planning the primary and supply network at a regional level requires a different approch that integrates elements of policy-making, standard engineering practice and considerable investments. This part of the system is extremely critical as failures here can propogate and lead to loss of service for thousands of people and businesses. Beyond system reliability and deep uncertainties, the complex governance requirements, such as the coordination between national, regional, and local administrations, present a major obstacle to integrated policy-making and long-term planning of these systems.
Over the past decade, the field has steadily refined and built upon several approaches that address WDS planning under uncertainties. To begin, the introduction of staged design, where interventions are modeled in phases rather then in as single batch representing the target network, provided a tool to represent long-term plans with more realism (for an example on WDS see Creaco et al., 2014). Next, this modeling framework was extended to include uncertainty, either through a robust approach that accounts for uncertainty (e.g., Creaco et al., 2015), or through a flexible approach where one plan with multiple pathways (in its simplest form: a decision tree) is devised, but only one branch is implemented based on the uncertainties realisation (e.g., Basupi and Kapelan, 2015).
An inherent limitation of robust and flexible plans is that they require accounting for all possible future system interventions and scenarios at the initial planning stage. Therefore, when uncertainties move away from the accounted values or new opportunities emerge (i.e., under deep uncertainties), these approaches fall short: the water utilities must adapt to the changes in their environment.
An adaptive plan would typically feature at least one of two components. The first is a periodic re-evaluation and adjustment of the plan based on uncertainty evolution to incorporate new opportunities and information (see, for an example, Beh et al., (2017) on water supply systems, and Skerker et al., (2023) on “the value of learning” in flexible planning for water reservoirs). The second is the ability to shift between pathways and avoiding locking-up to one (see dynamic adaptive policy pathways for the Rhine Delta, Haasnoot et al., 2013).
Regardless of definitions and classifications, the field also lacks coherence in benchmark definition, as none of these approaches have been systematically tested against each other. At best, works introducing different methodologies may have used the same benchmark networks (e.g., Anytown or New York Tunnels), but the overall problem descriptions—and thus the assumptions and scenarios surrounding these networks—differ substantially.
Therefore, in the context of the Battle of the Water Futures, the participants’ goal is to apply novel methodologies to develop masterplans for a nation’s water utilities under deep uncertainties; while the orgainsers objective is to provide a benchmark that advances the state of practice in decision-making under deep uncertainty for water distribution systems and establishes a foundation for benchmarking through a highly customizable, open-source framework.
1.3 Ambition
Participants must develop masterplans for a nation’s water utilities operating across different provinces. However, the BWF presents a level of detail and computational complexity such that the featured transport network optimization problem cannot be solved through brute force optimization. Moreover, unlike traditional optimization problems with specified scenarios, here, only a description of the environment dynamics and a handful of expert-driven scenarios for key drivers (e.g., population) are provided. Competitors must embrace uncertainty and decide for themselves what to model and how, reflecting the reality that different teams would make different modeling choices when facing deep uncertainty.
The decision to place participants in a situation of deep uncertainty is deliberate. While some parameters are uncertain and the evolution of these uncertainty bounds in the future will not be disclosed (e.g., the annual cost of operating a production facility, which varies between minimum and maximum values), past observations will be made available at the end of each competition stage, enabling participants to implement adaptive strategies and learn from accumulated evidence. Similarly, the competition ranking process has not been announced, only the system requirements are explained in this document (i.e., the metrics under which the system is evaluated), compelling participants to identify their own compromise solutions within this many-objective context.
2 Problem Description
The Battle of the Water Futures takes place in a not so ficticious world, with a blend of real data and assumptions. The BWF is concerned with the optimal planning and operation of multiple regional water distribution systems in a national landscape. It is based on The Netherlands where KWR is located and instead of the current water ulitities division, one independent water utility is assumed for every province. Each water utility will need to prepare a masterplan for at least the next 25 years and run the operations of their systems. Competitors will also make some decision regarding national and regional policies, like budget allocation and price increase. Various information about the current status of the system (network model, prices, demands, population, etc) are made available, consult DATA AVAILABLE page. Together wit hthe current status, future projections of the relevant drivers (e.g., population, electtricity shares) can be used to formulate the plan. The masterplan consist of a series of interventions, policy decisions and operations that can be applied over the years of the planning horizon, this effectively requires staged design and applied phased interventions.
Once the masterplan is completed, the utilities will start following it and operate and mantain the system accordingly to the specifications. The world will evolve, uncertainties will be realized (shown) and, hopefully, they will be able to fullfill their primary objective: satisfying the consumers water demand with adequate level of service.
During the time when the masterplan is being executed, several key performance indicator are extracted, from the main finance health and reliability (level of service), to secondary objectives like GHG emissions and fairness. The water utilities will also have access to data about their networks, such as pressures, demands, power consumptions etc
After 25 years, the utilities will have a chance to update their masterplan, adapting it based on the new acquired informaiton (now history), and the cycle will repeat (plan, execute, observe).
Both interventions and input parameters can be assumed to be deep uncertaintites, meaning that only approximate definition of their future evolution are provided and that no probability (or wrong ones) are associated with the realisation of each scenario. Everything should be regarded only as an indication, to the best of the current knowledge, made by some experts. Take these as an example: - the time to construct a desalination plant: it can be defined between 5 and 10 years. Once construction starts its actual implementation time can be anything between those bounds - population evolution: projections of population in the country wille be provided with a mean value and upper and lower bounds with various uncertainty level. However, how this parameter will evolve no one can know Behind the scenes we blend experts opinions with a dynamic model with tunable parameters to to generate the synthetic time serieses that will be the input of this system.
Planning horizon: free, but at least 25 years Interventions resolution (masterplan intervnetions timestep): year/quarter Stage horizon/masterplan update period: 25 years Hydraulic simulation timestep: 1 hour Every other parameter will have its own resolution as defined in the input files. Simulation approach: every day for the whole stage > if too computationally demanding (>20 minute per solution), we reduce the number of days mantaining the same timestep, i.e., 1 week representative of each month or 1 week representative of each quarter/season.
Definitions Deep uncertainties Staged Design Robust planning Flexible planning Adpative planning
2.1 System Decription
This chapter describes the physical components that constitute the water transport network in the BWF competition. We present the key entities—municipalities, water sources, pumping stations, and pipes—detailing their characteristics, operational constraints, and how they evolve over time. Each component’s properties, costs, and decision variables are outlined to provide participants with a complete understanding of the system they will manage. The physical feasibility of any network configuration is verified through hydraulic simulation in EPANET. External factors affecting the system, such as climate conditions and the electricity grid, are addressed separately in Section 2.2.
2.1.1 Water Utilities
Within the BWF, Water Utilities (WUs) act as the entities responsible for managing the assets and delivering drinking water across their assigned provinces.
WUs must chose the policies to apply to their own system (e.g, setting the budget for non-revenue water reduction) and decide the interventions to take on their system (e.g., replacing a pipe).
Water utilities can also join their efforts to pursue a common interventions, for example, the cost of placing a pipe connecting two water utilities will be shared between the two utilities.
The structure of the system is static. The number of WUs and their specific geographic responsabilities (assigned provinces) will not change through the competition.
The key parameters and decision variables governing the Water Utility Module are detailed in Table LABEL:tbl:wu-properties. The actual values for these variables can be inspected within the data files, which are mapped in Appendix A.
Table 1: Water utilities’ properties review.
Property Type Scope
Identifier Static Water Utility
Assigned province Static Water Utility
2.1.2 Municipalities
Each municipality (gemeente) is represented as a single junction point with positive demand, abstracting the entire secondary and tertiary distribution network within that jurisdiction. This node serves as the sole supply point for all water within the municipality’s boundaries, with characteristics such as population, land area, and housing stock consolidated at this level.
The system presents two main challenges. First, municipal parameters evolve over time as cities grow and change. Second, the network topology itself is dynamic: municipalities can merge or be absorbed by larger neighbors, causing the number of nodes to vary throughout the planning horizon.
Excursus on the Modelling Approach
To model this administrative restructuring, municipalities can only open or close on January 1st of each year. When a municipality closes, its delivery point disappears from the network.
Absorption by existing municipalities: When a municipality is absorbed by a larger neighbor that already exists, all attributes of the closing municipality (population, land area, housing stock, etc.) transfer to the destination municipality. Any pipe that previously connected these two entities becomes hidden, as it formally becomes part of the destination municipality’s internal distribution network.
Clustering into new municipalities: When multiple municipalities close and cluster together to form a new entity, all their delivery points disappear and a new supply point emerges at the location of the newly formed municipality. The new municipality inherits all pipeline connections and attributes from the closing municipalities. This modelling approach mirrors real-world dynamics in densely populated countries like the Netherlands. When a new municipality forms through clustering, typically a new city center is established while former city centers become secondary neighborhoods. These moments of urban reorganization present natural opportunities for water utilities to lay new connections and redesign substantial portions of the distribution system.
Municipalities have many attributes that influence the other modules of the system. The full list can be seen in Table LABEL:tbl:muni-properties, while the actual values for these variables can be inspected within the data files, which are mapped in Appendix A.
Table 2: Municipalities’ properties review.
Property Type Scope Unit
Name Static Municipality
Identifier Static Municipality
Latitude Static Municipality degrees
Longitude Static Municipality degrees
Elevation Static Municipality m
Province Static Municipality
Begin date Static Municipality date
End date Static [Optional] Municipality date
End reason and destination Static [Optional] Municipality
Population Dynamic Exogenous Municipality inhabitants
Surface land Dynamic Exogenous Municipality $km^{2}$
Surface water (inland) Dynamic Exogenous Municipality $km^{2}$
Surface water (open water) Dynamic Exogenous Municipality $km^{2}$
Number of houses Dynamic Exogenous Municipality units
Average age distribution network Dynamic Endogenous Municipality years
Along with these attributes, the only observable quantity for each municipality is the total water demand divided in its two components: consumption (delivered demand) and the quota of undelivered demand.
More precisely, the total water demand comprises two volumetric quantities, though this breakdown is not observable to participants:
  • billable water demand (the sum of household and business water demands) described in Section 2.1.2.1, and
  • non-revenue water (accounting for leaks, flushing, measurement errors and other losses), described in Section 2.1.2.2.
Instead, only two outputs are observable: actual consumption and undelivered demand. These two variables are extracted from an EPANET simulation of the network run in pressure-driven analysis (PDA) mode with a minimum pressure threshold of 30 m. This threshold accounts for the requirement to deliver 20 m of pressure at the consumer tap, plus an allowance of 5 m pressure drop in the secondary network and 5 m in the tertiary portion.
2.1.2.1 Water Demand Model
The methodology developed to generate water consumption time series builds on historical data from the Dutch Drinking Water Statistics(Vewin 2025) , which provide nationwide trends in total drinking water production, sectoral water use, and non-revenue water over the period 2000–2024. Specifically, water-consumption time series generation is structured into three phases.
Phase I. The first phase estimates the annual water volume supplied to each municipality using information on households, and businesses, complemented by projected data where required. These annual volumes are calibrated to match national totals reported in official statistics and then randomized around the calibrated value to introduce variability among municipalities.
Phase II. In the second phase, representative hourly consumption profiles are assigned to each municipality using a library of year-long, normalized profiles derived from district-metered areas and pre-processed to remove leakage effects. In greater detail, for each municipality, two residential profiles are selected from the library according to municipality population class, while a single non-residential profile is drawn from a dedicated set.
Phase III. The third phase produces the final hourly time series by applying a Fourier series-based approach which combines seasonal modulation, climate-related adjustments, and random perturbations to capture temporal variability. The two residential profiles associated with each municipality are aggregated through weighted combinations, and both residential and non-residential profiles are scaled to match the previously estimated yearly volumes. The outcome is a set of 8,760-point hourly water consumption series for each municipality.
Table 3: Water demand model’s properties review.
Property Type Scope Unit
Population Dynamic Exogenous Municipality inhabitants
Number of houses Dynamic Exogenous Municipality units
Number of businesses Dynamic Exogenous Municipality units
Daily per household demand Dynamic Exogenous Municipality L/house/day
Daily per business demand Dynamic Exogenous Municipality L/business/day
Max yearly temperature Dynamic Exogenous National [°C]
2.1.2.2 Non-Revenue Water Model
Our modelling of non-revenue water (NRW) relies on the average age of the pipe infrastructure in a municipality. Based on this average age, the municipality gets assigned to one of five possible classes (A - E) as follows:
Age in years Class Lower and upper bound of NRW demand in mˆ3/day
0 - 25 A (0, 12)
25 - 43 B (12, 20)
43 - 54 C (20, 35)
54 - 60 D (35, 55)
> 60 E (55, inf)
Each class is associated with a distribution of non-revenue water (NRW)demands (e.g., leakages), from which we sample to generate the total demand, consisting of the normal demand + NRW demand. Notably, the NRW demand is different for each class – i.e., older systems suffer from more leaks and therefore have a higher NRW demand. The distribution of NRW demands per class and km of pipes is illustrated in the following figure.
Figure 1: leak-demands
The total number of km of pipes in a given municipality is linked to its population size. Here, we use a linear relationship between the population size and the km of pipes, also illustrated in the following Figure:
km_pipes=57.7population_in_10kkm\_pipes=57.7*population\_in\_10k
Figure 2: pop-pipes
Table 5: Non-revenue water model’s properties review.
Property Type Scope
NRW demand Static [Uncertain] National
Inner distribution network - length to population ratio Static National
Inner distribution network - length Dynamic endogenous Municipality
Inner distribution network - average age Dynamic endogenous Municipality
NRW intervention - unit cost Dynamic endogenous NRWClass, Municipality Size Class, National
NRW intervention - effectiveness factor Static [Uncertain] NRWClass, Municipality Size Class, National
Intervention budget Option
Intervention policy Option
2.1.3 Water Sources
Water sources represent the entire process of water abstraction and treatment consolidated at a single location. Each source has a nominal production capacity which represents a hard limit on the amount of water that can be delivered in any 24-hour window.
In the BWF, three macro-types of water sources are modelled: groundwater, surface water, and desalination. The main trade-off is that groundwater sources are generally smaller (< 30 $Mm^{3}/year$) and cheaper to operate, as the water has higher quality and requires less treatment. Surface water treatment plants and desalination plants are generally much larger (30-60 $Mm^{3}/year$) but are considerably more expensive and energy-intensive to run. Moreover, surface water sources differ as they are affected by climate conditions; specifically, low inflows within rivers could temporarily shut down treatment plants downstream (effectively reducing their capacity to zero for those days). Instead, groundwater sources have an extraction permit (expressed in $m^{3}$ per year). This is not a hard physical constraint such as the nominal production capacity, but rather a “soft” legislative constraint checked by the government at the end of each year. This penalty effectively represents compensation for the hydrological displacement affecting farmers and natural areas due to the overextraction. The fine amount is set by law and can therefore change at any time based on political decisions.
The cost of water production at every source is a combination of four components:
  • Fixed costs (personnel, taxes, planned maintenance, etc.)
  • Volumetric costs for energy
  • Volumetric costs for non-energy related expenditures (chemicals, filters, etc.)
  • Extra volumetric costs for non-energy related expenditures This final component is an additional cost paid when the source operates above its planned production capacity, i.e., nominal capacity multiplied by its capacity target factor (which can be interpreted as the ideal efficiency point).
Similarly to municipalities, sources can also open and close over time. Participants can decide to close production locations and open new ones within the constraints of the problem (available locations and sizes) to make the supply system more efficient. However, a closed source cannot be reopened and no direct cost is associated with this action. When activating a new source, participants must decide the nominal capacity, but the possible size is limited by different rules depending on the source type:
  • Desalination and surface water sources have an upper limit defined in the competition data (see Appendix A)
  • Groundwater sources cannot exceed the permit by more than 30%
New water sources have an uncertain construction time, so participants must communicate the construction start date, and the activation date will be randomized. The capital investment needed for a new source is simply the unit cost of construction at that year by the required size.
One limitation of the BWF is that we do not model water quality differences between sources, which would typically prevent mixing in practice. This simplification keeps the problem tractable given its existing complexity.
The key parameters and decision variables governing the Water Sources Module are detailed in Table LABEL:tbl:sources-properties. The actual values for these variables can be inspected within the data files, which are mapped in Appendix A.
Table 6: Sources’ properties review.
Property Type Scope Unit
Name Static [Optional] Source
Identifier Static Source
Source type Static Source
Latitude Static Source degrees
Longitude Static Source degrees
Elevation Static Source m
Province Static Source
Connected municipality Static Source
Activation date Static Source date
Closure date Static [Optional] [Decision] Source date
Capacity - nominal Static [Optional] [Decision] Source $m^{3}/day$
Capacity - target factor Static Source Type %
Unit cost of construction Dynamic Endogenous Source Size Class $\times$ National $\texteuro/m^{3}\dot{d}ay$
Operational costs - fixed Dynamic Endogenous [Uncertain] Source Size Class $\times$ National $\texteuro/year$
Operational costs - volumetric for energy Dynamic Endogenous [Uncertain] Source Size Class $\times$ National $kWh/m^{3}$
Operational costs - volumetric for non-energy Dynamic Endogenous [Uncertain] Source Size Class $\times$ National $\texteuro/m^{3}$
Operational costs - volumetric for non-energy - multiplier Static Source Type %
Construction time Static [Uncertain] Source Type years
Availability factor Dynamic Exogenous Surface water sources
Permit Static Groundwater sources $m^{3}/year$
Fine amount Dynamic Exogenous Groundwater source permit excedance Severity Class $m^{3}/year$
2.1.4 Pumping stations
Pumping stations are the infrastructure connecting water sources to municipalities and are responsible for distributing the water produced by each source into the network. Each pumping station contains one or more identical pumps operating in parallel.
Whenever a new source is opened, participants must define which pump option (selected exclusively from a pre-defined list of available pumps) and how many units are installed at that source’s pumping station. Similarly, participants can decide to replace pumps at existing locations.
While a water source’s daily outflow is constrained by its nominal capacity, the pumping station’s configuration limits the source’s peak outflow rate ($m^{3}/hour$).
The only operational cost component of a pumping station is its energy expenditure, calculated during the network hydraulic simulation based on the energy consumption and tariffs. Peak demand charges are not included, as we assume utilities have agreements in place with electrical grid providers (see more details in Section 2.2.2). Maintenance costs and other fixed yearly operational costs are included in the initial construction cost of each pump.
Pumps performances remain constant over time (no degradation in efficiency or similar wear effects). However, pumps do age normally and have an expected lifetime. The actual lifetime is randomized, and when a pump reaches its end of life, it must be replaced. Participants do not need to communicate this decision, as replacements will be automatically implemented, but they must account for this “unexpected” cost in their planning.
Table 7: Pumping stations’ properties review.
Property Type Scope Identifier Static Pumping station Assigned source Static Pumping station Pump option Static [Decision] Pumping station Number of pumps Static [Decision] Pumping station
Table 8: Pumps’ properties review.
Property Type Scope
Identifier Static Pump option
Pump curve Static Pump option
Efficiency curve Static Pump option
Expected lifetime Static [Uncertain] Pump option
Note that the provided pump curves only describe the working range of the pumps and must not be extrapolated beyond the values presented in the tables.
2.1.5 Connections (Pipes)
Within the BWF, there is a distinction between connections and pipes. Connections are the entities that define the possible links between the network’s nodes, while pipes are the actual physical elements installed to transport water. This distinction models the fact that different pipes can be installed on the same connection at different points in time. However, duplicate pipes are not allowed on the same connection.
Similarly to pumps, participants can install new pipes on any unused connection, or replace existing pipes on connections that already have one. Selection must occur from a predefined set of pipe options. Pipe options are characterized by material, hydraulic properties, installation cost, and associated carbon emissions, with a detailed overview available in Table LABEL:tbl:pi-properties and actual values provided in the data files mapped in Appendix A.
The Darcy friction factor of a new pipe is provided for every option. However, the rate at which the friction factor increases over time (decay rate) is uncertain and is bounded between minimum and maximum values.
Associated carbon emissions (in kg CO2-eq per meter of pipe installed) are also provided. These emission factors may change over time due to technological advancements.
Connections are either completely within a province (intra-province) or shared between provinces (inter-province). The complete lists of intra-province and inter-province connections are included in the data files mapped in Appendix A.
Table 9: Connections’ properties review.
Property Type Scope Unit Identifier Static Connection Node A Static Connection Node B Static Connection Type Static Connection Distance Static Connection m (Minor loss coefficient) set to 0
Table 10: Pipes’ properties review.
Property Type Scope Unit
Identifier Static Pipe option
Diameter Static Pipe option
Material Static Pipe option
Darcy friction factor - new pipe Static Pipe option
Darcy friction factor - decay rate Static [Uncertain] Pipe option
Darcy friction factor - existing pipe Dynamic endogenous Pipe option
Cost (new pipe) Dynamic endogenous Pipe option €/m
Equivalent emissions (new pipe) Dynamic exogenous Pipe option tCO2eq/m
2.2 System External Drivers
This chapter describes the external drivers that influence the water distribution system and how they have been modelled in the BWF competition. The key external systems modelled are: - the climate, - the energy system, and - the economic context.
These systems are beyond the direct control of the participants and introduce the main exogenous factors shaping the environment in which the transport network operates.
2.2.1 Climate & Hydro-meteorological Forcing
2.2.2 Energy System
We will model the following three different parameters for the evaluation of the water distribution networks.
Variable Variable Type Presented By
Elecricity price in €/kWh Dynamic Exogenous National (Simulation model)
Greenhouse gas emission by electricity production in kgCO2/kWh Dynamic Exogenous National (Simulation model)
Unit costs for solar pv panel in €/kW Dynamic Exogenous National (Simulation model)
We assume yearly changes in electricity price roughly influenced by inflation but with a high uncertainty. Due to different electricity demands during the day, the electricity price will change hourly following a daily pattern with peaks during the working hours and lows during the night. For each kWh obtained from the power suppliers we will estimate the produced greenhouse gas emissions from generating the energy. This value may change yearly due to different compositions of the electricity generation. The alternative to buying electricity from a power supplier is to generate it from solar pv panels that also come with specific costs. These costs may change yearly due to more efficiency and different climate conditions.
The action the participants can do is to install solar pv panels on pump stations at any time, that will have a lifespan of 25 years. The main objective is to achieve low electricity costs in the water distribution network. As a secondary objective the greenhouse gas emissions produced by the water distribution network should be minimized.
2.2.3 Economy
2.2.3.1 Bond Issuance
The bond issuance governs how additional capital is raised when a utility exhausts its allocated budget for a period. When the available budget is fully utilized, a bond is automatically issued to cover the costs of planned interventions. The price and yield of this bond are determined by the modeled bond demand and market conditions specific to the issuance year. These bonds introduce future repayment obligations, and all associated interest payments accumulate as part of the total financial cost of the interventions. While utilities cannot directly control bond yields, they can anticipate debt accumulation through scenario analysis and adopt strategies that maintain financial sustainability. For instance, utilities may design pacing of interventions to minimize borrowing, or evaluate alternative interventions that reduce the likelihood of large bond issuances during periods with unfavorable yield conditions.
Variable Variable Type Presented By
Remaining budget in year t Dynamic Endogenous Utility ID
Bond issuance trigger (budget exhausted) Dynamic Endogenous Utility ID
Bond amount issued Dynamic Endogenous Utility ID
Bond price (issue price) Dynamic Exogenous Simulation model
Bond yield Dynamic Exogenous Simulation model
Bond demand curve Dynamic Exogenous Simulation model
Interest payments Dynamic Endogenous Utility ID
Total financial cost (capex + interest) Dynamic Endogenous Utility ID
2.2.3.2 Budget Allocation
The budget allocation module governs how national funding is distributed across provincial water utilities, each of which operates and maintains its own water distribution system. This allocation affects the resources available to implement interventions and support operations. Budget distribution can follow several principled rules: proportional allocation to population size; inverse-proportional schemes favoring smaller regions; income-based allocations reflecting socioeconomic considerations; inverse-proportional schemes favoring less wealthy regions; or fully customized schemes.
Variable Variable Type Presented By
National total budget Static National
Population size Dynamic Endogenous Municipality ID
Income index (socio-economic) Dynamic Endogenous Municipality ID
Per-capita allocation weight Option Rule
Inverse-population allocation weight Option Rule
Income-based allocation weight Option Rule
Equity priority score (favoring less wealthy regions) Option Rule
Customized allocation vector Option Participant-defined
2.2.3.3 Inflation Dynamics
Inflation represents the annual rate at which general price levels evolve over the planning horizon. In the BWF setting, inflation affects all monetary flows: capital expenditures, operating expenses, various costs, and prices for energy consumption and water tariffs. Because inflation compounds over time, it can significantly alter the real purchasing power of fixed budgets and shape the financial feasibility of long-term interventions. Interventions on inflation are limited, as inflation is an exogenous macroeconomic indicator outside the direct control of water utilities. However, utilities may adopt strategies that account for inflation, such as front-loading major investments to avoid future cost escalation, or stress-testing the interventions under alternative inflation trajectories. The primary goal is to design actions and financial flows that remain robust despite uncertain future price conditions.
Variable Variable Type Presented By
Inflation rate in year t Dynamic Exogenous National (Simulation model)
2.2.3.4 Water Tariffs
The water tariffs module models how utilities generate revenue through a pricing scheme commonly used in real-world systems: a fixed component (e.g., service charge) combined with a volumetric component (per-m³ charge). Both components evolve annually according to the prevailing inflation rate, ensuring that real revenue tracks changes in general price levels. Both the fixed and volumetric tariff components are exogenously given and evolve solely according to the uncertain inflation dynamics defined in the scenario generator. Participants must therefore design plans that remain financially robust under tariff paths they cannot control. This constraint emphasizes the importance of planning under deep uncertainty: utilities must cope with potential revenue volatility driven by inflation.
Variable Variable Type Presented By
Fixed tariff component (service charge) Dynamic Exogenous Utility ID
Volumetric tariff component [€/m³] Dynamic Exogenous Utility ID
Annual tariff escalation rule Dynamic Exogenous National (inflation-indexed)
2.3 System Interventions
This chapter provides a summary of all possible interventions available to participants in the BWF competition. All interventions have been introduced in previous sections. Here, we consolidate this information to give participants a complete overview of the decision space they can navigate when developing their strategies.
2.3.1 Non-Revenue Water Interventions
Participants can decide how much (as a percentage) of each water utility’s yearly budget should be spent on improving pipe infrastructure to reduce non-revenue water. Improving the pipe infrastructure means decreasing its average age, which is linked to the different leak classes and implies the amount of non-revenue water. There exist two policies that distribute a utility’s budget across its municipalities:
  • by_leak_class Uses the budget to improve the leak class (by one) of each municipality in a greedy way (worst cases first) until no budget is left.
  • by_population Distributes the budget proportionally to the municipalities’ population size.
2.4 System Requirements
This chapter describes the metrics used to evaluate each solution (masterplan) and the overall system performance in the BWF competition. These metrics define the many-objective space that participants must navigate, covering the main aspects of economic performance, environmental impact, reliability, and fairness. Importantly, these are evaluation metrics rather than explicit objectives: the competition ranking formula will not be disclosed to participants. The ranking may use weighted aggregation (with weights potentially changing over time to reflect shifting social priorities) or custom multi-criteria evaluation methods (e.g., selecting only solutions above a reliability threshold before ranking by other criteria). However, participants may choose to treat these metrics as objectives in their own optimization processes when developing their masterplans.
2.4.1 Total Annualized Cost
  • Formula: $TAC=\sum_{j}\frac{K_{j}}{L_{j}}+OPEX+(coupon×P_{issue})$
  • Calculated over all dimensions (utility, time). Ranked based on cumulative amount across time and utilities.
  • unit: €
2.4.2 Total Annualized GHG Emissions
  • Formula: $\text{GHG}=\sum_{j}\frac{K_{j}\times EF^{\text{emb}}_{j}}{L_{j}}+\sum_{t}\frac{E_{t}\times EF}{1000}$
  • Calculated over all dimensions (utility, time). Ranked based on cumulative amount across time and utilities.
  • unit: tCO2eq
2.4.3 Service Reliability
  • Formula: $R^{\text{service}}=1-\dfrac{\sum_{i,t}U_{i,t}}{\sum_{i,t}D_{i,t}}$
  • Calculated over all dimensions (municipality, household class, and time). Participants will be ranked based on one specific combination of these dimensions, which is kept uncertain and can change between stages.
  • unit: dimensionless between 0 and 1 (higher is better)
2.4.4 Affordability Fairness
  • Formula: $AF=\frac{p_{v}×V_{lifeline}+F_{fixed}}{I_{p20}}×100$
  • Calculated over all dimensions (municipality, household class, and time). Participants will be ranked based on one specific combination of these dimensions, which is kept uncertain and can change between stages.
  • unit: % (lower is better)
This section defines all variables and notation used in the KPI formulas.
Symbol Units Description
General / Indices
i Index of node or municipality
t Time index (e.g., hourly step)
j Index of intervention
N - Total number of nodes
T hours Total number of simulation hours in the evaluation period
Demand Variables
$D_{i,t}$ m³/h Water demand at node i and time t
Qi,tQ{i,t} m³/h Delivered volume at node i and time t
$U_{i,t}=\max(D_{i,t}-Q_{i,t},0)$ m³/h Undelivered demand at node $i$, time $t$
Financial Variables
$K_{j}$ Capital expenditure (CAPEX) for intervention j
$L_{j}$ yr Asset lifetime of intervention j
OPEX €/yr Operating expenditures (energy, maintenance, etc)
$r_{f}$ Risk-free rate (e.g., 0.03 = 3 %)
cs Base credit spread (e.g., 0.01 = 1 %)
a Spread sensitivity to demand (e.g., 0.02 = 2 %)
d Random investor demand factor $\in$ [0.8, 1.2]
coupon Coupon rate of issued bond = $r_{f}+cs+a(1–d)$
$P_{\text{issue}}$ Principal amount of issued bond
Environmental Variables
$E_{t}$ kWh Pumping energy at time t
EFEF kg CO2/kWh Electricity emission factor
$EF^{emb}_{j}$ t CO2 / € or unit Embedded emission factor for intervention j
Socio-Economic Variables
$p_{v}$ €/m³ Volumetric tariff
$F_{\text{fixed}}$ €/month Fixed charge per household
$V_{\text{lifeline}}$ m³/month Lifeline volume (e.g., 1.5 m³/person/month)
$I_{p20}$ €/month 20th-percentile monthly income across nodes
Acknowledgements
Thanks to everyone.
Data Availability Statement
Reference to the Zenodo repo.
Mention the data providers and the licenses.
[…]
Here we will describe all the file in the zip folder and which variables they have.