### v4

# The Battle of the Water Futures

The Battle of the Water Futures
Dennis Zanutto    Christos Michalopoulos    Lydia Tsiami    André Artelt    Jasmin Brandt    Demetrios Eliades    Stelios Vrachimis    Stefano Alvisi    Valentina Marsili    Filippo Mazzoni    Panagiotis Samartzis    Barbara Hammer    Phoebe Koundouri    Marios Polycarpou    Dragan Savić
(November 2025)
Executive Summary
This competition challenges participants to confront one of water distribution systems’s most pressing problems: adaptive planning under deep uncertainty. Participants will develop a 75-year national masterplan for a hypothetical Dutch water grid that must balance four competing objectives: financial sustainability, service reliability, environmental impact, and equitable access. Unlike traditional planning exercises, this competition requires strategic decisions across three 25-year stages while navigating uncertain climate, energy, and economic scenarios. Success demands developing an adaptive masterplan that blends policy decisions and infrastructure interventions, gets updated based on evidence accumulated across the implementation stages, and preserves flexibility to avoid locking into suboptimal pathways. This document defines the competition framework; a subsequent report will present participant solutions and outcomes.
1 Introduction
1.1 Disclaimer
The networks and the problem presented in this competition are synthetic and, despite being based on the Netherlands, do not represent any real-world situation. While we utilize real data from trusted sources where possible (e.g., CBS [^1] CBS (Centraal Bureau voor de Statistiek) is the Dutch national statistics agency: https://www.cbs.nl for municipal data, KNMI [^2] KNMI (Koninklijk Nederlands Meteorologisch Instituut) is the Dutch meteorological institute: https://www.knmi.nl for weather observations, Vewin [^3] Vewin is the association of Dutch drinking water companies: https://www.vewin.nl for statistics about water utilities), other aspects rely on assumptions based on expert knowledge.
We are committed to the model’s integrity and, based on competitor feedback, we will continuously test and refine the model (e.g., balancing electricity and infrastructure costs) to ensure it remains physically sound and balanced. If you identify any significant inconsistencies, please inform us and we will review them promptly.
All versions are archived on Zenodo, with the latest version establishing the official standard (Zanutto et al. 2025). An open-source evaluator will also be released in the WaterBenchmarkHub (Artelt et al. 2025), allowing competitors to test their solutions before the final submission.
1.2 Background and motivation
Water distribution systems (WDS) grow organically with cities and are therefore inherently subject to deep uncertainties. Strategic WDS planning focuses on the long-term design of the primary supply system – the national grid. While secondary and tertiary network components can typically be selected using established engineering principles, planning the primary and supply network at a regional level requires a different approach, one that integrates policy considerations, standard engineering practice, and substantial investment. This part of the system is extremely critical. Failures at this level can propagate and lead to loss of service for thousands of people and businesses. Beyond system reliability and deep uncertainties, the complex governance requirements, such as the coordination between national, regional, and local administrations, present a major obstacle to integrated policy-making and long-term planning of these systems.
Over the past decade, the field has steadily refined and built upon several approaches that address WDS planning under uncertainties. To begin, the introduction of staged design, where interventions are modelled in phases rather than in a single batch representing the target network, provided a tool to represent masterplans with more realism (e.g., Creaco, Franchini, and Walski 2014). Next, this modelling framework was extended to include uncertainty, either through a robust approach that accounts for it (e.g., Creaco, Franchini, and Walski 2015), or through a flexible approach that develops a single plan with multiple pathways. In its simplest form, this takes the shape of a decision tree, where only one branch is implemented based on how uncertainties unfold (e.g., Basupi and Kapelan 2015).
An inherent limitation of robust and flexible plans is that they require accounting for all possible future system interventions and scenarios at the initial planning stage. Therefore, when uncertainties move away from the expected values or new opportunities emerge (i.e., under deep uncertainties), these approaches fall short, and the water utilities must adapt to the changes in their environment.
An adaptive plan would typically feature at least one of two components. The first component involves periodically re-evaluating and adjusting the plan as uncertainties evolve, allowing new opportunities and information to be incorporated. For example, see Beh, Maier, and Dandy (2015) on water supply systems, and Skerker et al. (2023) on “the value of learning” in flexible planning for water reservoirs. The second is the ability to shift between pathways and avoid locking up to one (see dynamic adaptive policy pathways for the Rhine Delta, Haasnoot et al. 2013).
Regardless of definitions and classifications, the field also lacks coherence in benchmark definition, as none of these approaches have been systematically tested against each other. At best, works introducing different methodologies may have used the same benchmark networks (e.g., Anytown or New York Tunnels), but the overall problem descriptions—and thus the assumptions and scenarios surrounding these networks—differ substantially.
Therefore, in the context of this Battle of the Water Futures (BWF), the participants’ goal is to apply innovative methodologies to develop a masterplan for a national water grid under deep uncertainties. The organisers, in turn, seek to establish a benchmark that advances decision-making practices for water distribution systems under deep uncertainty, while laying the ground for future benchmarking through a highly customizable, open-source framework.
1.3 Ambition
Participants must develop masterplans for a national water grid with water utilities operating across different provinces. However, the BWF presents a level of detail and computational complexity such that the featured national grid optimisation problem cannot be solved through brute force optimisation. Moreover, unlike traditional optimisation problems with specified scenarios, here, only a description of the environment dynamics and a handful of expert-driven scenarios for key drivers (e.g., population) are provided. Competitors must embrace uncertainty and decide for themselves what to model and how to do so, reflecting the reality that different teams would make different modelling choices when facing deep uncertainty.
The decision to place participants in a context of deep uncertainty is intentional. The BWF will unfold over three competition stages, each spanning 25 years, to realistically emulate the long-term challenges faced by planners. This means that some parameters are uncertain and their uncertainty bounds evolution will be provided only through approximate expert-based scenarios (e.g., population projections vary between minimum and maximum values, though actual outcomes may fall outside these bounds). At the end of each competition stage, past observations will be made available, enabling participants to implement adaptive strategies and learn from accumulated evidence. Similarly, the competition ranking process has been deliberately left undisclosed; only the system requirements, namely, the metrics under which the system is evaluated, are outlined in this document. This approach is more akin to realistic situations and compels participants to identify their own compromise solutions within this many-objective context.
2 Problem Description
The Battle of the Water Futures (BWF) problem focuses on a hypothetical national drinking water grid of the Netherlands. The system structure (demand nodes, sources, and links), available interventions (e.g., pipes, treatment plants), and exogenous drivers (e.g., population, climate) are based on publicly available data about the Dutch context, supplemented by assumptions or insights from other regions as needed.
The main objective of the BWF is to keep the system operational until the end of the century, thereby preventing bankruptcy while fulfilling the required service requirements in each water utility. Secondary objectives include reducing emissions toward climate neutrality and ensuring fairness across regions and generations. To achieve these goals, the country’s water utilities come together to develop a national masterplan, which they agree to update every 25 years.
The masterplan specifies the interventions to be applied each year, both at the national level (e.g., connecting independent water utilities) and at the individual utility level (e.g., opening new sources). The masterplan does not need to include the replacement of ageing components (pipes and pumps), as these are automatically replaced when they reach their end of life. However, to manage this uncertainty, utilities can either schedule replacements in advance or allow components to age naturally and account for replacement costs in the budget. The masterplan also defines the national and regional policies and whether they are revised during the planning period (note: while interventions must be specified annually, policies can be set initially and remain as they are until amended in the plan).
To prepare the masterplan, information on the current system status is made available, including network models, prices, demands, population, and other relevant parameters. Future projections of key drivers (e.g., population growth, climate trends) are also provided to complement system knowledge. However, these scenarios are based on current expert knowledge and should not be treated as precise forecasts as these variables are generally deep uncertainties of the system.
Examples of deep uncertainties include:
  • Desalination plant construction time: Planners are confident it can be built in 5 years, but acknowledge that delays could extend completion to 10 years.
  • Population evolution: Population is expected to follow the mean values or remain within the upper and lower bounds described by the national statistics agency. However, major external events, such as conflicts or mass migration driven by climate catastrophes, could push evolution beyond these ranges.
Behind the scenes, expert opinions are combined with dynamic models to generate the synthetic time series that drive system evolution. At each timestep, EPANET hydraulic simulations verify the network’s physical capability to transfer water and meet demands under the proposed interventions and operating conditions.
Once the masterplan is defined, utilities begin following it, operating and maintaining their systems according to its specifications. Around them, the world evolves and uncertainties unfold. After 25 years, utilities reconvene to update their plan. All system information is updated, and the performance over the past 25 years along with observed variables can be used to revise or completely redraw the masterplan. This cycle—plan, execute, observe, adapt—repeats for three rounds, testing the utilities’ ability to manage both short-term operations and long-term strategic planning under deep uncertainty.
2.1 System Description
This chapter describes the physical components that constitute the water transport network in the BWF competition. The key entities are municipalities, water sources, pumping stations, and pipes and each of the following section will detail their characteristics, operational constraints, and how they evolve over time. The physical feasibility of any network configuration is verified through hydraulic simulation in EPANET. External factors affecting the system, such as climate conditions and the electricity grid, are addressed separately in Section 2.2.
2.1.1 Water Utilities
Within the BWF, Water Utilities (WUs) act as the entities responsible for managing the assets and delivering drinking water across their assigned provinces.
WUs must choose the policies to apply to their own system (e.g, setting the budget for non-revenue water reduction) and decide the interventions to take on their system (e.g., replacing a pipe).
Multiple WUs can also join their efforts to pursue common interventions, for example, the cost of placing a pipe connecting two water utilities will be shared between the two utilities. If WU are connected with each other and there is an exchange of water, the receiving utility pays a volumetric fee to the supplying utility. More detailed information about the water tariff scheme are available in Section 2.2.3.3.
The structure of the system is static. The number of WUs and their specific geographic responsibilities (assigned provinces) will not change through the competition.
The key parameters and decision variables governing the Water Utility Module are detailed in Table LABEL:tbl:wu-properties. The actual values for these variables can be inspected within the data files, which are mapped in Appendix A.
Table 1: Water utilities’ properties review.
Property Type Scope Unit
Identifier Static Water Utility
Assigned province Static Water Utility
2.1.2 Municipalities
Each municipality (gemeente) is represented as a single junction point with positive demand, abstracting the entire secondary and tertiary distribution network within that jurisdiction. This node serves as the sole supply point for all water within the municipality’s boundaries, with characteristics such as population, land area, and housing stock consolidated at this level.
The system presents two main challenges. First, municipal parameters evolve over time as cities grow and change. Second, the network topology itself is dynamic: municipalities can merge or be absorbed by larger neighbors, causing the number of nodes to vary throughout the planning horizon.
Excursus on the Modelling Approach
To model this administrative restructuring, municipalities can only open or close on January 1st of each year. When a municipality closes, its delivery point disappears from the network.
Absorption by existing municipalities: When a municipality is absorbed by a larger neighbor that already exists, all attributes of the closing municipality (population, land area, housing stock, etc.) transfer to the destination municipality. Any pipe that previously connected these two entities becomes hidden, as it formally becomes part of the destination municipality’s internal distribution network.
Clustering into new municipalities: When multiple municipalities close and cluster together to form a new entity, all their delivery points disappear and a new supply point emerges at the location of the newly formed municipality. The new municipality inherits all pipeline connections and attributes from the closing municipalities. This modelling approach mirrors real-world dynamics in densely populated countries like the Netherlands. When a new municipality forms through clustering, typically a new city center is established while former city centers become secondary neighborhoods. These moments of urban reorganization present natural opportunities for water utilities to lay new connections and redesign substantial portions of the distribution system.
Municipalities have many attributes that influence the other modules of the system. The full list can be seen in Table LABEL:tbl:muni-properties, while the actual values for these variables can be inspected within the data files, which are mapped in Appendix A.
Table 2: Municipalities’ properties review.
Property Type Scope Unit
Name Static Municipality
Identifier Static Municipality
Latitude Static Municipality degrees
Longitude Static Municipality degrees
Elevation Static Municipality m
Province Static Municipality
Begin date Static Municipality date
End date Static [Optional] Municipality date
End reason and destination Static [Optional] Municipality
Population Dynamic Exogenous Municipality inhabitants
Surface land Dynamic Exogenous Municipality $km^{2}$
Surface water (inland) Dynamic Exogenous Municipality $km^{2}$
Surface water (open water) Dynamic Exogenous Municipality $km^{2}$
Number of houses Dynamic Exogenous Municipality units
Number of businesses Dynamic Exogenous Municipality units
Average Disposable Income Dynamic Exogenous Municipality $k\text{\texteuro}$
Average age distribution network Dynamic Endogenous Municipality years
The total municipality water demand comprises two volumetric quantities:
  • Billable water demand ($D^{\text{BIL}}$): The sum of household and business water demands described in Section $2.1.$
  • Non-revenue water ($D^{\text{NRW}}$): accounting for leaks, flushing, measurement errors and other losses described in Section 2.1.2.2.
While these quantities represent the physical components of demand, they are not directly observable by participants. Instead, participants observe the total water demand divided into two components: consumption ($Q$; delivered outflow) and undelivered demand ($U$).
For each municipality $m$ at time $t$, these quantities maintain the following relationships:
$D_{m}(t)$ $=D^{\text{BIL}}_{m}(t)+D^{\text{NRW}}_{m}(t)$ (1)
$=Q_{m}(t)+U_{m}(t)$
The decomposition between delivered and undelivered demand is extracted from an EPANET simulation of the network run in pressure-driven analysis (PDA) mode with a minimum pressure threshold of 30 m. Whenever there is undelivered demand, we assume that this reduces the billable component first, i.e., $Q^{\text{BIL}}_{m}(t)=D^{\text{BIL}}_{m}(t)-U_{m}(t)$.
2.1.2.1 Water Demand Model
The methodology developed to generate water consumption time series builds on historical data from the Dutch association of water companies (Vewin 2025), which provide nationwide trends in total drinking water production, sectoral water use, and non-revenue water over the period 2000–2024. Specifically, water-consumption time series generation is structured into three phases.
Phase I. The first phase estimates the annual water volume supplied to each municipality using information on households and businesses (CBS 2025), complemented by projected data where required. These annual volumes are calibrated to match national totals reported in official statistics (Vewin 2025) and then randomized around the calibrated value to introduce variability among municipalities.
Phase II. In the second phase, representative hourly consumption profiles are assigned to each municipality using a library of year-long, normalized profiles derived from district-metered areas and pre-processed to remove leakage effects. In greater detail, for each municipality, two residential profiles are selected from the library according to municipality population class, while a single non-residential profile is drawn from a dedicated set.
Phase III. The third phase produces the final hourly time series by applying a Fourier series-based approach which combines seasonal modulation, climate-related adjustments (accounting for the maximum yearly temperature), and random perturbations to capture temporal variability. The two residential profiles associated with each municipality are aggregated through weighted combinations, and both residential and non-residential profiles are scaled to match the previously estimated yearly volumes.
Therefore, the total billable demand of municipality $m$ at time $t$ (within year $y$) is defined as:
[…]
where $D^{\text{R1}}_{m}(t,T_{y})$ and $D^{\text{R2}}_{m}(t,T_{y})$ represent the two residential demands, $w_{m}\in[0,1]$ is the unitary weight to combine them, $D^{\text{C}}_{m}(t,T_{y})$ is the (commercial) non-residential demand, and $T_{y}$ is the maximum temperature recorded in year $y$.
Table 3: Water demand model’s properties review.
Property Type Scope Unit
Population Dynamic Exogenous Municipality inhabitants
Number of houses Dynamic Exogenous Municipality units
Number of businesses Dynamic Exogenous Municipality units
Daily per household demand Dynamic Exogenous Municipality m³/house/hour
Daily per business demand Dynamic Exogenous Municipality m³/business/hour
Max yearly temperature Dynamic Exogenous National °C
2.1.2.2 Non-Revenue Water Model
Non-revenue water (NRW) is an uncertain quantity modeled through the average age of pipe infrastructure in each municipality’s inner distribution network (IDN). Based on this average age, municipalities are assigned to one of five NRW classes as reported in Table LABEL:tbl:nrw-classes. Each class is associated with a distinct probability distribution of NRW demands, from which daily samples are drawn to generate the volumetric NRW demand factor ($m^{3}/km/day$). Notably, older infrastructure suffers from more leaks and therefore exhibits higher NRW demand factors.
The distribution of NRW demands varies by class and is illustrated in Figure 1. The actual values of the distributions’ parameters can be inspected within the data files, which are mapped in Appendix A.
The total length of pipes in a municipality is linked to its population size through the following linear relationship:
[…]
where $m$ is the municipality index, $y$ is the year, and $L^{\text{IDN}}_{m}(y)$ is the total length of pipes (km) in municipality $m$ at year $y$.
The actual municipality NRW demand is also capped at twice the billable daily demand to prevent unrealistic leakage levels. Therefore, total NRW demand for municipality $m$ at day $d$ is calculated as:
[…]
where $f^{\text{NRW}}_{\text{class}(m)}$ is the sampled NRW demand factor ($m^{3}/km/day$) for the municipality’s class, and $\bar{D}^{\text{BIL}}_{m}(d)$ is the average daily water demand of municipality $m$ at day $d$. The daily leak $D^{\text{NRW}}_{m}(d)$ is equally spread across the day (i.e., $D^{\text{NRW}}_{m}(t)=D^{\text{NRW}}_{m}(d)/24$).
Figure 1: Non-revenue water demand factor per class
Table 4: Non-revenue water classification by infrastructure age.
Inner distribution network - average age [years] NRW Class Probability distribution
0 - 25 A Inverted Exponential
25 - 43 B Uniform
43 - 54 C Uniform
54 - 60 D Uniform
> 60 E Exponential
Table 5: Non-revenue water model’s properties review.
Property Type Scope Unit
Inner distribution network - length to population ratio Static National $km\cdot(10^{4}\text{ inhabitants})^{-1}$
Inner distribution network - length Dynamic endogenous Municipality kmkm
Inner distribution network - average age Dynamic endogenous Municipality yearsyears
NRW intervention - unit cost Dynamic endogenous NRWClass, Municipality Size Class, National $\text{\texteuro}/year/km$
NRW intervention - effectiveness factor Static [Uncertain] NRWClass, Municipality Size Class, National
Intervention budget Option $\text{\texteuro}/year$
Intervention policy Option
2.1.3 Water Sources
Water sources represent the entire process of water abstraction and treatment consolidated at a single location. Each source has a nominal production capacity which represents a hard limit on the amount of water that can be delivered throughout any consecutive 24-hour window.
In the BWF, three macro-types of water sources are modelled: groundwater, surface water, and desalination. The main trade-off is that groundwater sources are generally smaller (< 25 $Mm^{3}/year$) and cheaper to operate, as the water has higher quality and requires less treatment. Surface water treatment plants and desalination plants are generally much larger (30-45 $Mm^{3}/year$) but are considerably more expensive and energy-intensive to run. Moreover, surface water sources differ as they are affected by climate conditions; specifically, low inflows within rivers could temporarily shut down treatment plants downstream (effectively reducing their capacity to zero for those days). Instead, groundwater sources have an extraction permit (expressed in $m^{3}$ per year). This is not a hard physical constraint such as the nominal production capacity, but rather a “soft” legislative constraint checked by the government at the end of each year. This penalty effectively represents compensation for the hydrological displacement affecting farmers and natural areas due to the overextraction. The fine amount is set by law and can therefore change at any time based on political decisions.
The cost of water production at every source $s$ is a combination of four components:
[…]
where:
  • $F_{s}$ represents fixed costs, including personnel, taxes, and planned maintenance.
  • $c^{\text{el}}(y)\cdot E^{\text{grid}}_{s}(y)$ represents volumetric costs for energy.
  • $c^{\text{vol}}_{s}(y)\cdot Q_{s}(y)$ represents the volumetric costs for non-energy related expenditures, such as chemicals and filters.
  • $c^{\text{extra}}_{s}(y)\cdot\max\bigl(Q_{s}(y)-\phi_{s}\cdot Q^{*}_{s},0\bigr)$ represent the extra volumetric costs incurred when production exceeds the planned threshold.
In Equation 5, $Q_{s}(y)$ is the total volume produced by the source in year $y$, $Q^{*}_{s}$ is the source nominal capacity, and $\phi_{s}$ is the capacity target factor (source-type dependent), which defines the ideal operating point above which additional costs are applied.
Additionally, the variable $E^{\text{grid}}_{s}(y)$ represents the electrical grid energy consumed for water production by the source $s$ in year $y$ and it is calculated as:
[…]
where $t$ is the simulation timestep, $\mathcal{Y}$ is the set of timesteps in year $y$, $E_{s}(t)$ represents the energy consumed by the sources $s$ at time $t$, and the $E^{\text{PV}}_{s}(t)$ is electricity generated by the photovoltaic panels associated with the source $s$ at time $t$.
The source energy consumption $E_{s}(t)$ is:
[…]
where $\epsilon_{s}$ is the source specific energy and $Q_{s}(t)$ is the volume produced by the source at timestep $t$.
To complete the accounting, the total operational expenditure associated with all the sources in water utility $w$ at year $y$ is:
[…]
where $\mathcal{S}_{w}$ represents the collection of water sources managed by utility $w$, and $\mathrm{OPEX}_{s}(y)$ is the individual source cost defined in Equation 5.
Similarly to municipalities, sources can also open and close over time. Participants can decide to close production locations and open new ones within the constraints of the problem (available locations and sizes) to make the supply system more efficient. However, a closed source cannot be reopened and no direct cost is associated with this action. When activating a new source, participants must decide the nominal capacity, but the possible size is limited by different rules depending on the source type:
  • Desalination and surface water sources have an upper limit defined in the competition data (see Appendix A)
  • Groundwater sources cannot exceed the permit by more than 30%
New water sources have an uncertain construction time, so participants must communicate the construction start date, and the activation date will be randomized.
The capital investment associated with the construction of new sources in water utility $w$ at year $y$ is:
[…]
where for a source $s$ in the set of the water utility’s sources ($\mathcal{S}_{w}$), $\tau_{s}$ is its starting construction time, $\mathbf{1}_{\{\tau_{s}=y\}}$ is an indicator function equal to 1 if the construction begins in year $y$ (0 otherwise), $Q^{*}_{s}$ is the requested source nominal capacity, and $c^{\text{source}}(\text{class}(s),y)$ the unit cost, which depends on the year of construction and the source class (see Table LABEL:tbl:sources-classes).
Table 6: Sources classification by their nominal capacity.
Source nominal capacity $Mm^{3}/year$] Source Class 4 SMALL 8 MEDIUM 16 LARGE > 16 VERY LARGE
One limitation of the BWF is that we do not model water quality differences between sources, which would typically prevent mixing in practice. This simplification keeps the problem tractable given its existing complexity.
The key parameters and decision variables governing the Water Sources Module are detailed in Table LABEL:tbl:sources-properties. The actual values for these variables can be inspected within the data files, which are mapped in Appendix A.
Table 7: Sources’ properties review.
Property Type Scope Unit
Name Static [Optional] Source
Identifier Static Source
Source type Static Source
Latitude Static Source degrees
Longitude Static Source degrees
Elevation Static Source m
Province Static Source
Connected municipality Static Source
Activation date Static Source date
Closure date Static [Optional] [Decision] Source date
Capacity - nominal Static [Optional] [Decision] Source $m^{3}/day$
Capacity - target factor Static Source Type %
Unit cost of construction Dynamic Endogenous Source Size Class $\times$ National $\text{\texteuro}/(m^{3}/day)$
Operational costs - fixed Dynamic Endogenous [Uncertain] Source Size Class $\times$ National $\text{\texteuro}/year$
Specific energy Static [Uncertain] Source $kWh/m^{3}$
Operational costs - volumetric for non-energy Dynamic Endogenous [Uncertain] Source Size Class $\times$ National $\text{\texteuro}/m^{3}$
Operational costs - volumetric for non-energy - multiplier Static Source Type %
Construction time Static [Uncertain] Source Type years
Availability factor Dynamic Exogenous Surface water sources
Permit Static Groundwater sources $m^{3}/year$
Fine amount Dynamic Exogenous Groundwater source permit excedance Severity Class
2.1.4 Pumping stations
Pumping stations are the infrastructure connecting water sources to municipalities and are responsible for distributing the water produced by each source into the network. Each pumping station contains one or more identical pumps operating in parallel.
Whenever a new source is opened, participants must define which pump option (selected exclusively from a pre-defined list of available pumps) and how many units are installed at that source’s pumping station. Similarly, participants can decide to replace pumps at existing locations or add new ones of the same type as installed.
Whether replacing old pumps or being added in new pumping stations, the capital investment associated with the installation of pumps in water utility $w$ at year $y$ is:
[…]
where $n$ is the pumping station index within the set of the water utility’s pumping stations ($\mathcal{N}_{w}$), $p$ is the p-th pump in that pumping station set ($\mathcal{P}_{n}$), $\tau_{p}$ is the pump installation time, $\mathbf{1}_{\{\tau_{p}=y\}}$ is an indicator function equal to 1 if the installation happened in year $y$ (0 otherwise), and $c^{\text{pump}}(\gamma_{p},y)$ the unit cost for the selected pump option $\gamma_{p}$.
While a water source’s daily outflow is constrained by its nominal capacity, the pumping station’s configuration limits the source’s peak outflow rate ($m^{3}/hour$).
The only operational cost component of a pumping station is its energy expenditure, calculated during the network hydraulic simulation based on the energy consumption and tariffs. Peak demand charges are not included, as we assume utilities have agreements in place with electrical grid providers (see more details in Section 2.2.2). Maintenance costs and other fixed yearly operational costs are included in the initial construction cost of each pump.
Thus, the operational expenditure associated with pumping in water utility $w$ at year $y$ is:
[…]
where $n$ is the pumping station index within the set of the water utility’s pumping stations ($\mathcal{N}_{w}$), $c^{\text{el}}(y)$ is the dynamic electricity price, and $E^{\text{grid}}_{n}(y)$ is the electricity consumption from the grid for the n-th pumping station in year $y$. This quantity is computed as:
[…]
where $t$ is the simulation timestep, $\mathcal{Y}$ is the set of timesteps in year $y$, $E_{p}(t)$ the p-th pump energy consumption at time $t$, $\mathcal{P}_{n}$ the set of pumps in pumping station $n$, and $E^{\text{PV}}_{n}(t)$ is electricity generated by the photovoltaic panels associated with pumping station $n$ at time $t$.
Pumps performances remain constant over time (no degradation in efficiency or similar wear effects). However, pumps do age normally and have an expected lifetime. The actual lifetime is randomized, and when a pump reaches its end of life, it must be replaced. Participants do not need to communicate this decision, as replacements will be automatically implemented, but they must account for this “unexpected” cost in their planning.
Table 8: Pumping stations’ properties review.
Property Type Scope Unit Identifier Static Pumping station Assigned source Static Pumping station Pump option Decision Pumping station Number of pumps Decision Pumping station unit
Table 9: Pumps’ properties review.
Property Type Scope Unit
Identifier Static Pump option
Pump curve Static Pump option
Efficiency curve Static Pump option
Lifetime Static [Uncertain] Pump option years
Unit cost Dynamic Endogenous Pump option
Note that the provided pump curves only describe the working range of the pumps and must not be extrapolated beyond the values presented in the tables.
2.1.5 Connections (Pipes)
Within the BWF, there is a distinction between connections and pipes. Connections are the entities that define the possible links between the network’s nodes, while pipes are the actual physical elements installed to transport water. This distinction models the fact that different pipes can be installed on the same connection at different points in time. However, duplicate pipes are not allowed on the same connection.
Similarly to pumps, participants can install new pipes on any unused connection, or replace existing pipes on connections that already have one. Selection must occur from a predefined set of pipe options. Pipe options are characterized by material, hydraulic properties, installation cost, and associated carbon emissions, with a detailed overview available in Table LABEL:tbl:pi-properties and actual values provided in the data files mapped in Appendix A.
Whether replacing old pipes or being added in a connection, the capital investment associated with the installation of pipes in water utility $w$ at year $y$ is:
[…]
where $\mathcal{J}_{w}$ is the set of all pipes within water utility $w$, $j$ the pipe index, $\tau_{j}$ is the pipe installation time, $\mathbf{1}_{\{\tau_{j}=y\}}$ is an indicator function equal to 1 if the installation happened in year $y$ (0 otherwise), $c^{\text{pipe}}(\theta_{j},y)$ the unit cost for the selected pipe option $\theta_{j}$, and $L_{j}$ is the length of pipe $j$.
The Darcy friction factor of a new pipe is provided for every option. However, the rate at which the friction factor increases over time (decay rate) is uncertain and is bounded between minimum and maximum values.
Associated carbon emissions (in kg CO2-eq per meter of pipe installed) are also provided. These emission factors may change over time due to technological advancements.
Connections are either completely within a province (intra-province) or shared between provinces (inter-province). The complete lists of intra-province and inter-province connections are included in the data files mapped in Appendix A. The capital cost for inter-provincial pipes is shared equally between the water utilities operating in the connected provinces (i.e., Equation 13 is used but the pipe’s length is half).
Table 10: Connections’ properties review.
Property Type Scope Unit Identifier Static Connection Node A Static Connection Node B Static Connection Type Static Connection Distance Static Connection m Pipe option Decision Connection Pipe installation date Decision Connection (Minor loss coefficient) set to 0
Table 11: Pipes’ properties review.
Property Type Scope Unit
Identifier Static Pipe option
Diameter Static Pipe option
Material Static Pipe option
Darcy friction factor - new pipe Static Pipe option
Darcy friction factor - decay rate Static [Uncertain] Pipe option $years^{-}1$
Darcy friction factor - existing pipe Dynamic endogenous Pipe option
Lifetime Static [Uncertain] Pipe option yearsyears
Unit cost (new pipe) Dynamic endogenous Pipe option $\text{\texteuro}/m$
Equivalent emissions (new pipe) Dynamic exogenous Pipe option $\text{tCO2eq}/m$
2.2 System External Drivers
This chapter describes the external drivers that influence the national WDS and how they have been modelled in the BWF competition. The key external systems modelled are the climate, the energy system, and the economic context.
These systems are beyond the direct control of the participants and introduce the main exogenous factors shaping the environment in which the transport network operates.
2.2.1 Climate & Hydro-meteorological Forcing
Climate is an exogenous factor beyond the control of the national water sector. Four meteorological variables are relevant for the national grid planning: temperature, precipitation, solar radiation and evotranspiration.
The average maximum temperature affects water demand seasonality.
Temperature, precipitation, and evotranspiration influence river flows (Meuse and Rhine). When these flows fall below mandated environmental thresholds, surface water treatment plants must reduce operations or close, temporarily limiting available supply.
All four variables affect groundwater availability and recharge rates, potentially reducing the permitted extraction volumes of new sources in the future. For a more detailed description of surface water and groundwater source dynamics, see Section 2.1.3.
Solar radiation influences the long-term electricity generation potential of solar panels, as sustained higher irradiance levels lead to greater cumulative energy yields over time.
The review of the climate variables can be seen in Table LABEL:tbl:clim-properties, while the actual values for these variables can be inspected within the data files, which are mapped in Appendix A. Note that instead of the evotranspiration variable, we directly provide the Standardized Precipitation-Evapotranspiration Index (SPEI) as it helps to quantify the severity, duration, and frequency of droughts by indicating whether conditions are wetter or drier than normal over a specific period.
Table 12: Climate model’s properties review.
Property Type Scope Unit
Temperature Dynamic Exogenous National °C
Precipitation Dynamic Exogenous National mm/day
Solar Radiation Dynamic Exogenous National W/m²
Stand. Precip.-Evap. Ind. (SPEI) Dynamic Exogenous National -
2.2.2 Energy System
The energy system in the BWF is modeled through three macro-drivers: electricity prices, greenhouse gas emission factor, and solar panels prices.
The electricity price is assumed to vary annually, influenced by inflation but subject to high uncertainty. On a higher resolution, prices also fluctuate hourly throughout the day. The electricity purchased by the grid has an associated greenhouse gas emissions factor, which can not be influcenced by the utilities but is adjusted yearly to account for the evolution of the national electricity generation mix.
As an alternative to grid electricity, participants may install solar photovoltaic panels. The unit cost of solar energy is modeled as a dynamic variable, changing annually to reflect improvements in the technology.
The capital investment associated with the installation of solar panels in water utility $w$ at year $y$ is:
[…]
where $v$ is the solar installation index within the set of photovoltaic systems in the water utility ($\mathcal{V}_{w}$), $\tau_{v}$ is the installation time, $\mathbf{1}_{\{\tau_{v}=y\}}$ is an indicator function equal to 1 if the installation happened in year $y$ (0 otherwise), $c^{\text{solar}}(y)$ is the unit cost per kilowatt of capacity in year $y$, and $P_{v}$ is the nominal power capacity of the solar system $v$.
No cost is associated with the operation of the solar panels. These components have a fixed lifespan of 25 years, after which competitors must choose between investing in a replacement installation or reverting to full dependence on the electrical grid.
Table 13: Energy system model’s properties review.
Property Type Scope Unit
Electricity price Dynamic Exogenous National $\text{\texteuro}/kWh$
Emission factor Dynamic Exogenous National $tCO2eq/kWh$
Unit cost - solar panel Dynamic Endogenous National $\text{\texteuro}/kW$
2.2.3 Economy
2.2.3.1 Inflation Dynamics
Inflation captures the year-over-year change in the general price level. In the BWF, it affects all costs equally except energy prices, which follow different dynamics see 2.2.2.
As inflation compounds over time, it can substantially erode water utilities’ purchasing power and alter the financial viability of capital-intensive interventions. While inflation is an exogenous macroeconomic variable beyond the direct control of water utilities, competitors can adopt strategic responses: front-loading major investments to lock in current prices, or stress-testing their masterplan against multiple inflation scenarios to ensure robustness.
In the absence of reliable long-term inflation forecasts, a reasonable baseline assumption is the central bank’s target rate (e.g., 2% for the European Central Bank).
Table 14: Inflation’s properties review.
Property Type Scope Unit
Inflation rate Dynamic Exogenous National
Note that wherever unit prices are given only in base-year terms (e.g., only the value in year 2000 is reported in the input files), their value in year $y$ ($C(y)$) is obtained by multiplying by the cumulative inflation index
[…]
where $C(0)$ is the base-year unit price and $\pi(y)$ is the inflation rate in year $y$.
2.2.3.2 Budget Allocation
A national budget (predefined over the planning horizon) is distributed annually across utilities to fund infrastructure interventions and operational expenses.
Competitors must decide how to distribute this budget among utilities in every year. Budget distribution can follow several principled rules: proportional allocation to population size; inverse-proportional schemes favuoring smaller regions; income-based allocations reflecting socioeconomic considerations; inverse-proportional schemes favuoring less wealthy regions; or fully customised schemes.
Table 15: Budget allocation model’s properties review.
Property Type Scope Unit
National total budget Static National €/year
Budget allocation policy Option Water utility
2.2.3.3 Water Pricing
Water utilities generate revenue selling water to three types of customers: residential users, businesses, and, possibly, other water utilities. Residential and commercial customers are billed using a two-part pricing scheme: a fixed service charge ($\text{\texteuro}/year$) and a volumetric charge based on consumption ($\text{\texteuro}/m^{3}$). For simplicity, both retail customer types face identical rates within each utility (with also no differentiation by income class or other categories). Water transactions between utilities, however, use only a volumetric charge based on the net exchange at the end of the year.
Thus, the total revenue for a water utility $w$ in year $y$ is:
[…]
where $P_{w}^{\text{fixed}}(y)$ and $P_{w}^{\text{variable}}(y)$ are the fixed and volumetric components of the retail water price, $Q^{\text{BIL}}_{m}$ is the delivered billable demand in municipality $m$, $\mathcal{M}_{w}$ is the set of municipalities served by water utility $w$, $P_{w}^{\text{sell}}(y)$ is the volumetric charge that utility $w$ applies for inter-utility water sales, $Q^{w^{\prime}+}_{w}(y)$ is the net positive volume of water sold by utility $w$ to utility $w^{\prime}$, and $\mathcal{W}^{-}$ is the set of water utilities excluding utility $w$.
The net water exchange between utilities $w$ and $w^{\prime}$ is defined as:
$\Delta Q_{w}^{w^{\prime}}(y)$ $=\sum_{t\in\mathcal{Y}}\sum_{j\in\mathcal{J}_{(w,w^{\prime})}}Q_{j}(t)$ (17)
$Q_{w}^{w^{\prime}+}(y)$ $=\max(\Delta Q_{w}^{w^{\prime}}(y),0)$
$Q_{w}^{w^{\prime}-}(y)$ $=-\min(\Delta Q_{w}^{w^{\prime}}(y),0)$
where $t$ is the simulation timestep, $\mathcal{Y}$ is the set of timesteps in year $y$, $Q_{j}(t)$ the flow over connection $j$ from the set of connections between the water utilities $\mathcal{J}_{(w,w^{\prime})}$ (with flow direction positive from $w$ to $w^{\prime}$).
Note that if a water utility has a negative net exchange with another utility (i.e., $\Delta Q_{w}^{w^{\prime}}(y)<0$), that will be regarded as a water import cost:
[…]
where $P_{w^{\prime}}^{\text{sell}}(y)$ is the volumetric charge applied by utility $w^{\prime}$ for water sales.
Participants must decide the water pricing adjustment strategy. They can either let all three quantities adjust according to inflation, or define a custom policy by specifying the yearly percentage increase for each of them independently.
Table 16: Water price model’s properties review.
Property Type Scope Unit
Fixed tariff component (service charge) Dynamic Endogenous Water utility €/year
Volumetric tariff component Dynamic Endogenous Water utility $\text{\texteuro}/m^{3}$
Water price for other utilities Dynamic Endogenous Water utility $\text{\texteuro}/m^{3}$
Water pricing adjustment policy Option Water utility
2.2.3.4 Bond Issuance
Whenever a water utility is unable to cover its expenditures in a given year, it finances the resulting deficit by issuing nationally backed bonds. Bonds are automatically generated to cover the utility debt in that year.
Specifically, the bond amount is $\kappa$ times the debt, i.e., $\mathrm{amount}_{i}=\kappa\cdot\mathrm{debt}_{w}(y)$ where $\kappa\in[1,2.5]$. A value of $\kappa$ equal to 1, implies that, depending on investor demand and prevailing market conditions, the proceeds from the bond issuance may barely cover the utility’s financing needs, leaving the balance for the following year around 0. In contrast, higher values of $\kappa$ (closer to 2.5) generate a cash surplus with the downside of a larger principal obligation to be repaid at maturity. Bonds are also characterised by a maturity of $M$ years, determining when the bond principal must be repaid, a coupon rate ($\mathrm{coupon}_{i}$), which determines the interest payments due each year, and a yield to maturity ($\mathrm{yield}_{i}$), which determines the price of the bond ($\mathrm{price}_{i}$).
Each year, the utility must repay the sum of principal amounts of all bonds reaching maturity plus the annual interest payments and and receives proceeds based on the bond price in that year. Formally:
$\text{PRI}_{w}(y)=\sum_{i\in\mathcal{B}_{w}(y):y=\tau_{i}+M}\mathrm{amount}_{i}$ (19)
$\text{INT}_{w}(y)=\sum_{i\in\mathcal{B}_{w}(y)}\mathrm{amount}_{i}\cdot\mathrm{coupon}_{i}$
[…]
where $i$ indicates the i-th bond, $\tau_{i}$ is the issuance year for bond $i$, $\mathcal{B}_{w}(y)$ is the set of bonds active for water utility $w$ in year $y$ and $\mathcal{B}_{w}(y):y=\tau_{i}+M$ the subset of bonds reaching maturity $M$.
The i-th bond’s coupon, yield, and price are given by:
$\mathrm{coupon}_{i}=r_{f}+\hat{\pi}(y=\tau_{i})$ (20)
$\mathrm{yield}_{i}=\mathrm{coupon}_{i}+a\cdot(1-d(y=\tau_{i}))$
$\mathrm{price}_{i}=\sum_{y=1}^{M}\frac{\mathrm{coupon}_{i}}{(1+\mathrm{yield}_{i})^{y}}+\frac{100}{(1+\mathrm{yield}_{i})^{M}}$
where $r_{f}$ is the risk-free rate (long-term government yield), $\hat{\pi}(y=\tau_{i})$ is the inflationary expectation at issuance year, $a$ is the sensitivity to investor demand, and $d(y=\tau_{i})$ is the uncertain demand factor for bond $i$ at issuance year.
Strong investor demand ($d(y)>1.0$) increases $\mathrm{PRO}_{w}(y)$ (through a lower yield and therefore higher price at issuance), while weak demand ($d(y)<1.0$) decreases it. This simulates real-world bond pricing where investor appetite introduces uncertainty to the utilities budgetting.
While utilities cannot directly control bond yields, they can anticipate debt accumulation through scenario analysis and adopt strategies that maintain financial sustainability. For instance, utilities may design pacing of interventions to minimize borrowing, or evaluate alternative interventions that reduce the likelihood of large bond issuances during periods with unfavorable yield conditions.
The complete list of the bond model properties can be seen in Table LABEL:tbl:eb-properties, while the actual values for these variables can be inspected within the data files, which are mapped in Appendix A.
Table 17: Bonds model’s properties review.
Property Type Scope Unit
Balance Dynamic Endogenous Water utility
Bond amount to debt ratio Option National
Bond amount Static Bond
Bond issue date Static Bond
Bond maturity Static National years
Risk free rate Static National
Inflationary expectations Dynamic Exogenous National
Bond coupon Dynamic Endogenous Bond
Sensitivity Static National
Investors demand factor Dynamic Exogenous National
Bond yield Dynamic Endogenous Bond
Bond price Dynamic Endogenous Bond
2.3 System Levers
Participants make two types of strategic decisions: policies and interventions.
This chapter provides a summary of all the options available to participants in the BWF competition (they have all been introduced in previous sections already).
Here, we consolidate this information to give participants a complete overview of the decision space they can navigate when developing their masterplan.
2.3.1 Policies
Policies encompass regulatory and operational rules, such as pricing structures, budget allocations, and maintenance protocols. Once set, policies remain in effect until explicitly amended.
2.3.1.1 National Budget Allocation (National)
Participants must decide the strategy to allocate the national budget across the water utilities. The policy can be a predefined one or follow a custom allocation of the funds:
  • by_population: Allocate the funds according to each province population.
  • by_inverse_population: Allocate the funds to provinces with less population (less revenue).
  • by_income: Allocate the funds according to each province economic activity.
  • by_inverse_income: Allocate the funds to provinces with less economic activity (less revenue).
  • custom: Allocate the budget according to the specified share for each water utility (must sum to 1).
year: 2025 national_policies: budget_allocation: policy: by_population # or "by_income" or the corresponding inverse policies
year: 2026 national_policies: budget_allocation: policy: custom policy_args: WU01: 0.12 WU02: 0.25 # and so on...
2.3.1.2 Non-Revenue Water Mitigation Budget (Utility)
Participants must decide each water utility’s yearly budget used to mitigate non-revenue water (NRW). This budget is used to improve the municipalities innner distribution network (IDN). More precisely, this budget descreases the municipalities IDN’s average age, which in turn improves the NRW class of the municipality, leading to a reduction of the NRW component. The policy can be a predefined one or follow a custom allocation of the funds:
  • by_nrw_class: Allocate the budget to improve by one NRW-class each municipality in a greedy way (worst cases first) until no budget is left in that year.
  • by_population: Allocate the budget according to each municipality population.
  • custom: Allocate the budget according to the specified share for each municipality (must sum to 1).
year: 2025 water_utility: WU01 policies: nrw_mitigation: budget: 30000 policy: by_nrw_class # or "by_population"
year: 2026 water_utility: WU01 policies: nrw_mitigation: budget: 30000 policy: custom policy_args: GM0001: 0.02 GM0002: 0.02 # and so on...
2.3.1.3 Water Pricing (Utility)
Participants must decide the water pricing adjustment strategy for each year. They have two options: increase all water price components according to inflation, or define a custom policy by specifying the percentage increase for each component (e.g., 2%).
year: 2025 water_utility: WU01 policies: pricing_adjustment: policy: by_inflation
year: 2026 water_utility: WU01 policies: pricing_adjustment: policy: custom # Specify the percentage increase for each pricing component policy_args: fixed_component: 0.03 # Annual increase for fixed costs (3%) variable_component: 0.02 # Annual increase for variable costs (2%) selling_price: 0.05 # Annual increase for water sales to other provinces (5%)
2.3.1.4 Bond Issuance (Utility)
Whenever the water utility is unable to cover its expenditures in a specific year, it finances the resulting deficit by issuing nationally backed bonds. Given that the raised amount is uncertain, participants can cover this uncertainty increasing the bond amount ratio. This adjustment is determined by the parameter $\kappa$, which can be any real number ranging from 1 to 2.5. More details are provided in Section 2.2.3.4.
year: 2025 water_utility: WU01 policies: bond_ratio: value: 2.0
2.3.2 Interventions
Interventions are physical modifications to the system, such as infrastructure upgrades or new installations. Interventions are specified annually and are always implemented. Note that if a utility contracts debt, a bond is issued automatically to cover the unbudgeted expenses.
2.3.2.1 Opening New Sources (Utility)
Participants can open new water sources to meet potential increases in demand. Available sources are predefined by location, and participants must specify a capacity within the allowable bounds.
year: 2025 water_utility: WU01 interventions: open_source: # Provide source identifier, capacity, and info about the connection - source_id: SG0158 source_capacity: 100 pump_option_id: PU003 n_pumps: 6 pipe_option_id: PI002 # Multiple sources can be added like this - source_id: SG0159 source_capacity: 50 pump_option_id: PU001 n_pumps: 4 pipe_option_id: PI003
2.3.2.2 Closing Sources (Utility)
Similarly, participants can close selected sources to improve the overall system efficiency.
year: 2025 water_utility: WU01 interventions: close_source: # Provide only the source identifier - source_id: SG0173 # Multiple sources can be removed like this. - source_id: SG0174
2.3.2.3 Installing Pipes (National or Utility)
Participants can decide to install new pipes or replace existing ones in the system. Each installation requires specifying the connection identifier and selecting a pipe option from a predefined list.
year: 2025 water_utility: WU01 interventions: install_pipe: # Provide connection identifier and pipe option - connection_id: CG0112 pipe_option_id: PI001 # Multiple pipes can be added like this. - connection_id: CG0113 pipe_option_id: PI008
2.3.2.4 Installing Pumps (Utility)
Participants can decide to install or replace pumps in pumping stations to calibrate the peak outflow of water sources. Each source is associated with one pumping station, which contains multiple identical pumps operating in parallel. Pump options must be selected from a predefined list. When installing pumps at an already-open pumping station, participants must specify whether to replace or add to the existing pumps. If the selected pump option differs from those already installed, all existing pumps will be automatically replaced, as only one pump type is allowed per station.
year: 2025 water_utility: WU01 interventions: install_pumps: # Provide source identifier, pump option and quantity for a new pumping station - source_id: SG0159 pump_option_id: PU003 n_pumps: 3 behaviour: replace # or "new"
2.3.2.5 Installing Solar (Utility)
Participants can decide to install behind-the-meter solar panels at water sources to reduce electricity costs and emissions for pumping stations. The solar panels offset electricity consumption and emissions but cannot be used as a profit-generating investment. Panels can be installed multiple times at different points in time. Note: Solar panels have a given lifespan (see Section 2.2.2); participants must decide whether to replace them upon expiration.
year: 2025 water_utility: WU01 interventions: install_solar: # Provide location and capacity - source_id: SG0158 capacity: 20 # Multiple sources can be added like this - source_id: SG0159 capacity: 20
2.4 System Requirements
This chapter describes the metrics used to evaluate each solution (masterplan) and the overall system performance in the BWF competition. These metrics define the many-objective space that participants must navigate, covering the main aspects of economic performance, environmental impact, reliability, and fairness. Importantly, these are evaluation metrics rather than explicit objectives: the competition ranking formula will not be disclosed to participants. The ranking may use weighted aggregation (with weights potentially changing over time to reflect shifting social priorities) or custom multi-criteria evaluation methods (e.g., selecting only solutions above a reliability threshold before ranking by other criteria). However, participants may choose to treat these metrics as objectives in their own optimization processes when developing their masterplans.
2.4.1 Maintain Financial Viability
Participants must seek to maintain all water utilities in a financially viable state throughout the planning horizon. While utilities should not pursue profit, they must avoid insolvency.
Each year ($y$), the financial balance $F_{w}$ of every water utility $w$ is updated based on the following set of equations:
$F^{*}_{w}(y+1)$ $=F_{w}(y)+\text{NIB}(y)\cdot\alpha_{w}(y)+\text{REV}_{w}(y)$ (21)
$\quad-\text{CAPEX}_{w}(y)-\text{OPEX}_{w}(y)$
$\quad-\text{WLR}_{w}(y)-\text{WIC}_{w}(y)$
$\quad-\text{INT}_{w}(y)-\text{PRI}_{w}(y)$
[…]
$F_{w}(y+1)$ $=F_{w}^{*}(y+1)+\text{PRO}_{w}(y)$
The provisional fund balance $F_{w}^{*}(y+1)$ is calculated by accounting for all inflows and outflows:
  • $\text{NIB}(y)\cdot\alpha_{w}(y)$ is the national investment budget allocated for the water utility $w$ (see Section 2.3.1.1),
  • $\text{REV}_{w}(y)$ is the water utility’s revenue from the billable water demand (Equation 16),
  • $\text{CAPEX}_{w}(y)$ represents all the utility’s interventions capital costs (i.e., the sum of Equations 9, 10, 13, 14),
  • $\text{OPEX}_{w}(y)$ accounts for all the utility’s operational costs (i.e., the sum of Equations 8, 11),
  • $\text{WLR}_{w}(y)$ is the budget for NRW mitigation (see Section 2.3.1.2),
  • $\text{WIC}_{w}(y)$ is the cost for imported water from other water utilities (Equation 18),
  • $\text{INT}_{w}(y)$ the utility’s interest payments due (Equation 19), and
  • $\text{PRI}_{w}(y)$ is the principal amount due (Equation 19).
If the provisional balance is negative, the deficit defines the debt ($\text{debt}_{w}(y)$), which triggers bond issuance. The bond proceeds $\text{PRO}_{w}(y)$ are determined according to Section 2.2.3.4 and Equations 19, 20.
The actual fund balance $F_{w}(y+1)$ is obtained by adding the bond proceeds (if any) to the provisional balance, ensuring the fund remains solvent. Note that while surpluses are carried forward, deficits are always financed through bond issuance automatically ensuring that the fund balance is always positive.
Therefore, evaluation will focus only on the remaining debt at the end of the planning period. Final debt measures insolvency risk, not financial optimality. Competitors are free to pursue any financial strategy they deem appropriate, such as minimising costs, adjusting prices to enable expensive solutions, or adopting other innovative approaches. The many-objective framework ensures balanced evaluation across all dimensions. In other words, maintaining financial stability is an achievable target; the real challenge is deciding what to sacrifice.
2.4.2 Minimize GHG Emissions
Participants must keep the masterplan Greenhouse Gas (GHG) emissions to the lowest feasible level.
The GHG emissions for water utility $w$ in year $y$ are:
[…]
where $\mathrm{GHG}_{w}^{\text{emb}}(y)$ the embedded (construction) emissions, and $\mathrm{GHG}_{w}^{\text{op}}(y)$ the operational emissions from electricity use.
Only new pipes have embedded GHG emissions. The embedded emissions in year $y$ are:
[…]
where $\mathcal{C}_{w}$ is the set of connections of water utility ww444as described in Section 2.1.5, we distinguish between connections and pipes, $\mathbf{1}_{c\text{ activated in }y}$ is 1 if connection $c$ installs a new pipe in year $y$, 0 otherwise, $EF_{p_{c}}(y)$ is the unit emission factor of the connection’s selected pipe option $p_{c}$, and $L_{c}$ is the connection length.
The unit emission factor $EF_{p}$ for pipe option $p$ depends on the pipe option diameter $Diam_{p}$ and material $Mat_{p}$ and may change over time because of technological advancements, i.e., $EF_{p}(y)=EF(Diam_{p},Mat_{p},y)$.
The operational emissions are calculated based on the total electricity purchased from the grid, covering both water treatment (sources) and transport (pumping).
[…]
where for each timestep $t$ of a year yy555yy represent the year, while $\mathcal{Y}$ is the collection of timesteps, $E_{s}(t)$ and $EF_{s}(t)$ are the energy consumption and the emission factor of source $s$, while $E_{p}(t)$ and $EF_{p}(t)$ represent the same quantities for pump $p$.
Pumps energy consumption is retrieved via the EPANET simulations, while the water sources energy consumption is calculated according to Equation 7\$.
The emission factors of both entities (pumping stations $EF_{p}(t)$ and sources $EF_{s}(t)$) are dynamic and depend on the size and time of production of the behind-the-meter solar panels installation at that location (if no solar is installed, this variable reduces to the constant electricity grid emission factor in year $y$, i.e., $EF^{\text{el}}(y)$ ).
2.4.3 Maximize Service Reliability
Participants must ensure high service reliability by minimizing unmet water demand.
Service reliability for municipality $m$ in year $y$ is:
[…]
where $U_{m}(y)$ is the undelivered demand and $D^{\text{BIL}}_{m}(y)$ is the billable water demand.
Evaluation will focus on maintaining adequate service levels across all municipalities of each water utility throughout the entire planning horizon.
2.4.4 Promote Affordability and Equity
Participants must ensure water remains affordable, particularly for lower-income households, while maintaining equitable pricing across municipalities.
The affordability fairness metric represents the fraction of income that a household at the 20th percentile of the income distribution would spend on essential water consumption. Affordability fairness (lower is better) for water utility $w$ in year $y$ is:
[…]
where $P_{w}^{\text{fixed}}(y)$ and $P_{w}^{\text{variable}}(y)$ are the fixed and variable components of water price, $D^{\text{life}}$ is lifeline volume (minimum water required per person), and $ADI_{w}^{p20}(y)$ is the 20th percentile of disposable income across all households served by the water utility.
Evaluation will focus on minimizing affordability while maintaining reasonable equity across utilities.
Acknowledgements
We thank all KWR colleagues for their invaluable suggestions and support throughout the project development. We would like to particularly thank Mirjam Blokker, Mark Morley, Karel van Laarhoven, Djordje Mitrovic, Ina Vertommen, Konstantinos Glynis, and Beatriz Gutierrez-Caloir.
Dennis, Lydia, Christos and Dragan would like to thank Yvonne Hassink-Mulder at Vitens for her insights on real water utility challenges; our collaboration profoundly shaped this battle from the early stages.
We would like to thank Luke Butler, whose tool EPANET.js was instrumental in developing this national network and enabled visualization of the information with unprecedented ease.
Dennis, Lydia and Christos would like to thank their academic supervisors Andrea Castelletti and Christos Makropoulos for their support and guidance.
D. Zanutto, C. Michalopoulos, L. Tsiami, J. Brandt, D. Eliades, S. Vrachimis, P. Samartzis, B. Hammer, P. Koundouri, M. Polycarpou, and D. Savić have received funding from the European Research Council (ERC) under the ERC Synergy Grant Water-Futures (Grant agreement No. 951424).
A. Artelt is supported by the Ministry of Culture and Science NRW (Germany) as part of the Lamarr Fellow Network.
This publication reflects the views of the authors only.
Data Availability Statement
All data versions are archived on Zenodo, with the latest version serving as the official standard (Zanutto et al. 2025). The files are also available in the releases section of the GitHub repository.
The model uses data from trusted open sources under Creative Commons licenses (e.g., CBS and KNMI) and draws on publicly available reports (e.g., Vewin). Where not explicitly cited, values are based on the authors’ institutional knowledge or represent modeling assumptions.
A complete description of all data sources will be released upon conclusion of the competition.
Notation
Entity Identifiers
Each entity is numbered sequentially starting from 1. Unique identifiers consist of two capital letters indicating the object type, followed by a number indicating the specific instance.
Jurisdictions:
  • Nation: NL0000
  • Region (Landsdeel): LD0000
  • Province: PV0000
  • Municipality (Gemeente): GM0000
Water Sources:
  • Groundwater: SG0000
  • Surface water: SS0000
  • Desalination: SD0000
Infrastructure:
  • Pumping stations: PS0000 (uniquely determines the source)
  • Pump option: PU000
Connections:
  • Between municipalities: CG0000
  • From source to municipality: CS0000
  • Between provinces: CP0000
  • Pipe option: PI000
Sets and Indices
[Our indexes here]
Parameters
[Our parameters here]
Decision Variables
[Our decision variables here]
[…]
This appendix describes the data contained within the supplementary material (zipped data folder). Four types of data are included in the folder and they are described independently in the following sections.
EPANET Networks Files
The provided network files were generated using EPANET 2.3. Unfortunately, most common EPANET libraries and software (e.g., WNTR and EPANET for Windows) are not keeping up with the versions and they may generate warnings upon opening or fail to open them at all. When using a library based on an outdated version of EPANET, the hydraulic results will also differ, as EPANET 2.3 features numerous bug fixes with respect to previous versions.
For these compatibility reason and to ensure consistency between results, the solutions will be submitted through an ad-hoc form and inp files will be rejected. The sole reference for correctness is the evaluator’s computations, which uses the latest EPANET 2.3 library.
At the time of writing, we suggest using EPANET.js to visualize the networks overlaid on a map of the Netherlands, while for hydraulic computations, we recommend using the EPANET-PLUS Python library, as it is currently the only one offering EPANET 2.3 support (Artelt 2025).
Configuration File
The configuration file configuration.yaml tells the evaluator how to build the system. Specifically, it indicates:
  • The relative paths of the Excel input files describing the system
  • The parameters and settings that control the simulation
This makes the configuration file the central reference point that connects all the input data sources with the simulation engine. By modifying these parameters (e.g., switching input files), participants can quickly simulate the system under different scenarios.
Settings
  • Start year (start_year) [ - ]
  • End year (end_year) [ - ]
  • National budget (national_budget) [€]
Evaluator Input Files
The evaluator input files fully describe the system entities (municipalities, sources, etc.) and modules (energy system, water demand module, etc.). By modifying these files, participants can simulate different scenarios and system configurations.
Static Properties Files
Static Properties files define the characteristics of each entity, with one entity per row and properties in columns. Multiple sheets can be used to distinguish between subtypes of the same object (e.g., water sources are divided into desalination, groundwater, and surface water).
Dynamic Properties Files
Dynamic Properties files contain time-varying data for both exogenous inputs and initial values of endogenous variables. Each property has its own sheet within the Excel file, organized as follows:
  • Rows: Each row represents a snapshot in time, with the first column containing the timestamp
  • Columns: Each column represents a scope, i.e., which entity or group of entities the value applies to. For example, while ‘GMxxxx’ indicates the value applies only to that specific municipality, ‘NL0000’ indicates a national scope and therefore applies to all municipalities.
When future values are reported (e.g., timestamp after 2024 for the first stage), they can be regarded as expert-driven scenarios. Actual values will override these parameters in subsequent stages of the competition.
Multi-Parameter (Static and Dynamic) Properties
When a property requires multiple parameters, column headers use a dash separator (‘-’) to indicate:
  • Uncertainty bounds: ['NL0000-min', 'NL0000-max']
  • Multi-dimensional scope: ['A-SMALL', 'A-MEDIUM', ...] (e.g., variable varying by both class and size)
  • Time series patterns: ['NL0000-1', 'NL0000-2', ...] (e.g., electricity price profiles at national level)
Modules
State
Entities
  • Entity: State
    • File: configuration.yaml
    • Properties:
        State name (name) [ - ]
        State ID (id) [ - ]
  • Entity: Region
    • File: jurisdictions/jurisdictions-static_properties.xlsx
    • Sheet: regions
    • Properties:
        Region name (name) [ - ]
        Region ID (cbs_id) [ - ]
        State ID (state) [ - ]
  • Entity: Province
    • File: jurisdictions/jurisdictions-static_properties.xlsx
    • Sheet: provinces
    • Properties:
        Province name (name) [ - ]
        Province ID (cbs_id) [ - ]
        Region ID (region) [ - ]
  • Entity: Municipality
    • File: jurisdictions/jurisdictions-static_properties.xlsx
    • Sheet: municipalities
    • Properties:
        Municipality name (name) [ - ]
        Municipality ID (cbs_id) [ - ]
        Province ID (province) [ - ]
        Opening date (begin_date) [ - ]
        Closure date (end_date) [ - ]
        Reason for the closure (end_reason) [ - ]
        Municipalities IDs that acquired the municipality’s assets (destination_cbs_ids) [ - ]
        Coordinates (latitude, longitude, elevation) [ degrees, degrees, m ]
        Area characteristics (touristic_area_cbs_id, touristic_group_cbs_id) [ -, - ]
Dynamic Properties
  • Property: Population
    • File: jurisdictions/municipalities-dynamic_properties.xlsx
    • Sheet: population
    • Scope: Municipality ID, National
    • Unit: [ inhabitants ]
    • Notes: National values are for future projection
  • Property: Surface land
    • File: jurisdictions/municipalities-dynamic_properties.xlsx
    • Sheet: surface-land
    • Scope: Municipality ID
    • Unit: [ $km^{2}$ ]
  • Property: Surface water (inland)
    • File: jurisdictions/municipalities-dynamic_properties.xlsx
    • Sheet: surface-water-inland
    • Scope: Municipality ID
    • Unit: [ $km^{2}$ ]
  • Property: Surface water (open water)
    • File: jurisdictions/municipalities-dynamic_properties.xlsx
    • Sheet: surface-water-open
    • Scope: Municipality ID
    • Unit: [ $km^{2}$ ]
  • Property: Number of houses
    • File: jurisdictions/municipalities-dynamic_properties.xlsx
    • Sheet: n_houses
    • Scope: Municipality ID
    • Unit: [ houses ]
  • Property: Number of businesses
    • File: jurisdictions/municipalities-dynamic_properties.xlsx
    • Sheet: n_businesses
    • Scope: Municipality ID
    • Unit: [ businesses ]
  • Property: Associated water demand pattern - Residential
    • File: jurisdictions/municipalities-dynamic_properties.xlsx
    • Sheet: assoc_dem_pat-residential
    • Dimension: Pair
    • Scope: Municipality ID
    • Unit: [ - ]
  • Property: Associated water demand pattern - Business
    • File: jurisdictions/municipalities-dynamic_properties.xlsx
    • Sheet: assoc_dem_pat-business
    • Scope: Municipality ID
    • Unit: [ - ]
  • Property: Average age of the Inner Distribution Network
    • File: jurisdictions/municipalities-dynamic_properties.xlsx
    • Sheet: dist_network-age-avg
    • Scope: Municipality ID
    • Unit: [ years ]
  • Property: Average disposable income
    • File: jurisdictions/municipalities-dynamic_properties.xlsx
    • Sheet: disposable_income-avg
    • Scope: Municipality ID
    • Unit: [ k€ ]
Water Demand Model
Entities
  • Entity: Water Demand Pattern
    • File: water_demand_model/water_demand_model-static_properties.xlsx
    • Sheet: residential, business
    • Properties:
        Water Demand Pattern ID (demand_pattern_id) [ - ]
        Pattern values (year_hour) [ - ] Dimension: hours of the year
Dynamic Properties
  • Property: Business Demand
    • File: water_demand_model/water_demand_model-dynamic_properties.xlsx
    • Sheet: per_business_demand
    • Scope: National
    • Dimension: Uniform Uncertain
    • Unit: [ $m^{3}/business/hour$ ]
  • Property: House Demand
    • File: water_demand_model/water_demand_model-dynamic_properties.xlsx
    • Sheet: per_house_demand
    • Scope: National
    • Dimension: Uniform Uncertain
    • Unit: [ $m^{3}/house/hour$ ]
Non-Revenue Water Model
Static Properties
  • Property: NRW Intervention success probability
    • File: configuration.yaml
    • Label: nrw_model-intervention_success_prob
    • Scope: National
    • Dimension: Uniform Uncertain
    • Unit: [ - ]
Dynamic Properties
  • Property: NRW Intervention Unit cost
    • File: jurisdictions/nrw_model-dynamic_properties.xlsx
    • Sheet: nrw_intervention-unit_cost
    • Scope: National
    • Dimension: NRW class $\times$ Municipality size class
    • Unit: [ $\text{\texteuro}/year/km$ ]
Sources
Entities
  • Entity: Water Source (Groundwater, Surface water, Desalination)
    • File: sources/sources-static_properties.xlsx
    • Sheet: groundwater, surface_water, desalination
    • Properties:
        Source Name (name) [ - ]
        Source ID (source_id) [ - ]
        Coordinates (latitude, longitude, elevation) [ degrees, degrees, m ]
        Province (province) [ - ]
        Closest municipality ID (closest_municipality)
        Activation date (activation_date) [ - ]
        Closure date (closure_date) [ - ]
        Nominal capacity (capacity-nominal) [ $m^{3}/day$ ]
        Specific energy (opex-volum-energy_factor) [ $kWh/m^{3}$ ]
[…]
        Basin (basin) [ - ] Only surface water
Static Properties
  • Property: Capacity Target Factor
    • File: sources/sources-static_properties.xlsx
    • Sheet: global
    • Label: capacity-target_factor
    • Scope: Source Type
    • Unit [ - ]
  • Property: Operational costs - volumetric for non-energy - multiplier
    • File: sources/sources-static_properties.xlsx
    • Sheet: global
    • Label opex-volum-other-multiplier
    • Scope: Source Type
    • Unit [ - ]
  • Property: Construction time
    • File: sources/sources-static_properties.xlsx
    • Sheet: global
    • Scope: Source Type
    • Dimension: Uniform Uncertain
    • Unit [ - ]
  • Property: Operational costs - volumetric for energy (specific energy)
    • File: sources/sources-static_properties.xlsx
    • Sheet: global
    • Scope: Source Type
    • Dimension: Uniform Uncertain
    • Unit [ - ]
    • Notes: sampled for new sources
Dynamic Properties
  • Property: Unit cost of construction
    • File: sources/{source_type}-dynamic_properties.xlsx
    • Sheet: new_source-unit_cost
    • Scope: National
    • Dimension: Source size
    • Unit: [ $\text{\texteuro}/(m^{3}/day)$ ]
  • Property: Operational costs - fixed (unit cost)
    • File: sources/{source_type}-dynamic_properties.xlsx
    • Sheet: opex-fixed
    • Scope: National
    • Dimension: Source size $\times$ Uniform Uncertain
    • Unit: [ $\text{\texteuro}/(m^{3})$ ]
    • Notes: gets multiplied by annual nominal capacity
  • Property: Operational costs - volumetric for non-energy
    • File: sources/{source_type}-dynamic_properties.xlsx
    • Sheet: opex-volum-other
    • Scope: National
    • Dimension: Source size $\times$ Uniform Uncertain
    • Unit: [ $\text{\texteuro}/(m^{3})$ ]
  • Property: Availability factor
    • File: sources/{source_type}-dynamic_properties.xlsx
    • Sheet: availability_factor
    • Scope: National, Basin
    • Unit: [ - ]
  • Property: Water displacement fine amount
    • File: sources/groundwater-dynamic_properties.xlsx
    • Sheet: water_displacement-fine_amount
    • Scope: National
    • Dimension: Displacement severity
    • Unit: [ € ]
Pumping Infrastructure
Entities
  • Entity: Pumping Station
    • File: pumping_stations/pumping_stations-static_properties.xlsx
    • Sheet: entities
    • Properties:
        Pumping Station ID (pumping_station_id) [ - ]
        Assigned source (assigned_source) [ - ]
        Installed pumps - options IDs (pumps-option_ids) [ - ]
        Installed pumps - installation date (pumps-installation_dates) [ - ]
        Installed pumps - decomission date (pumps-end_dates) [ - ]
  • Entity: Pump Options
    • File: pumps/pump_options-static_properties.xlsx
    • Sheet: options
    • Properties:
        Pump Option ID (option_id) [ - ]
        Name (name) [ - ]
        Nominal Flow rate (flow_rate-nominal) [ $m^{3}/hour$ ]
        Lifetime (lifetime) [ years ] Dimension: Uniform Uncertain
  • Entity: Pump Curve
    • File: pumps/pump_options-static_properties.xlsx
    • Sheet: {option_id}
    • Properties:
        Flowrate (flowrate) [ $m^{3}/hour$ ]
        Head (head) [ m ]
        Efficiency (efficiency) [ - ]
Dynamic Properties
  • Property: Unit cost for a new pump
    • File: pumps/pump_options-dynamic_properties.xlsx
    • Sheet: new_pump-cost
    • Scope: Pump Option
    • Unit: [ € ]
Piping Infrastructure
Entities
  • Entity: Connection
    • File: connections/connections-static_properties.xlsx
    • Sheet: provincial, sources, cross-provincial
    • Properties:
        Connection ID (connection_id) [ - ]
        From node (from_node) [ - ]
        To node (to_node) [ - ]
        Distance (distance) [ m ]
        Minor loss coefficient (minor_loss_coeff) [ - ]
        Installed pipes - options IDs (pipes-option_ids) [ - ]
        Installed pipes - installation date (pipes-installation_dates) [ - ]
  • Entity: Pipe Options
    • File: pipes/pipe_options-static_properties.xlsx
    • Sheet: options
    • Properties:
        Pipe Option ID (option_id) [ - ]
        Diameter (diameter) [ mm ]
        Material (material) [ - ]
        Darcy friction factor - New pipe (darcy_friction_factor-new_pipe) [ - ]
        Darcy friction factor - Decay rate (darcy_friction_factor-decay_rate) [ - ] Dimension: Uniform Uncertain
        Lifetime (lifetime) [ years ] Dimension: Uniform Uncertain
Dynamic Properties
  • Property: New Pipe Unit Cost
    • File: connections/pipe_options-dynamic_properties.xlsx
    • Sheet: new_pipe-unit_cost
    • Unit: [ $\text{\texteuro}/m$ ]
  • Property: New Pipe Emission Factor
    • File: connections/pipe_options-dynamic_properties.xlsx
    • Sheet: new_pipe-emissions_factor
    • Unit: [ $\text{tCO2eq}/m$ ]
Climate
Dynamic Properties
  • Property: Temperature (Average, Average of minimum/maximum daily temperatures, Maximum/Minimum temperature recorded)
    • File: climate/climate-dynamic_properties.xlsx
    • Sheet: temperature-avg, temperature-min-avg,temperature-max-avg, temperature-warmest_day, temperature-coldest_day
    • Scope: National
    • Unit: [ °C ]
  • Property: Precipitation
    • File: climate/climate-dynamic_properties.xlsx
    • Sheet: precipitation
    • Scope: National
    • Unit: [ $\text{mm}/day$ ]
  • Property: Solar Radiation
    • File: climate/climate-dynamic_properties.xlsx
    • Sheet: solar_radiation-avg
    • Scope: National
    • Unit: [ $\text{W}/m^{2}$ ]
  • Property: Standardized Precipitation-Evotranspiration Index
    • File: climate/climate-dynamic_properties.xlsx
    • Sheet: SPEI
    • Scope: National
    • Unit: [ - ]
Economy
Entities
  • Entity: Bonds
    • File: economy/bonds-static_properties.xlsx
    • Sheet: entities
    • Properties:
        Bond issuance ID (bond_issuance_id) [ - ]
        Number of bonds (n_bonds) [ - ]
        Issue date (issue_date) [ - ]
        Maturity date (maturity_date) [ - ]
        Coupon rate (coupon_rate) [ - ]
        Water Utility ID (water_utility_id) [ - ]
Dynamic Properties
  • Property: Inflation
    • File: economy/economy-dynamic_properties.xlsx
    • Sheet: inflation
    • Scope: National
    • Unit: [ % ]
  • Property: Inflation Expectation
    • File: economy/economy-dynamic_properties.xlsx
    • Sheet: inflation-expected
    • Scope: National
    • Unit: [ % ]
  • Property: Investor Demand
    • File: economy/economy-dynamic_properties.xlsx
    • Sheet: investor_demand
    • Scope: National
    • Unit: [ - ]
Energy System
Entities
  • Entity: Solar Farm
    • File: energy/solar_farms-static_properties.xlsx
    • Sheet: entities
    • Properties:
        Solar farm ID (solar_farm_id) [ - ]
        Capacity (capacity) [ - ]
        Installation date (installation_date) [ - ]
        Decommission date (decommission_date) [ - ]
        Connected entity ID source/pumping station (connected_entity_id) [ - ]
Dynamic Properties
  • Property: Electricity Price
    • File: energy/energy_system-dynamic_properties.xlsx
    • Sheet: electricity_price-unit_cost
    • Scope: National
    • Unit: [ $\text{\texteuro}/kWh$ ]
  • Property: Electricity Price Pattern
    • File: energy/energy_system-dynamic_properties.xlsx
    • Sheet: electricity_price-pattern
    • Scope: National
    • Dimension: Hour of the week
    • Unit: [ - ]
  • Property: Grid Emission factor
    • File: energy/energy_system-dynamic_properties.xlsx
    • Sheet: grid_emission_factor
    • Scope: National
    • Unit: [ $kgCO_{2}eq/kWh$ ]
  • Property: Solar Panel Unit Cost
    • File: energy/energy_system-dynamic_properties.xlsx
    • Sheet: solar_panel-unit_cost
    • Scope: National
    • Unit: [ $\text{\texteuro}/kW$ ]
Water Utilities
Entities
  • Entity: Water Utility
    • File: water_utilities/water_utilities-static_properties.xlsx
    • Sheet: entities
    • Properties:
        Identifier (water_utility_id) [ - ]
        Assigned provinces (assigned_provinces) [ - ]
Dynamic Properties
  • Property: Funds Balance
    • File: water_utilities/water_utilities-dynamic_properties.xlsx
    • Sheet: balance
    • Scope: Water Utility
    • Unit: [ € ]
  • Property: Water Price Fixed Component
    • File: water_utilities/water_utilities-dynamic_properties.xlsx
    • Sheet: water_price-fixed
    • Scope: Water Utility
    • Unit: [ €/house ]
  • Property: Water Price Variable Component
    • File: water_utilities/water_utilities-dynamic_properties.xlsx
    • Sheet: water_price-variable
    • Scope: Water Utility
    • Unit: [ $\text{\texteuro}/m^{3}$ ]
  • Property: Water Price Exchange Rate
    • File: water_utilities/water_utilities-dynamic_properties.xlsx
    • Sheet: water_price-selling
    • Scope: Water Utility
    • Unit: [ $\text{\texteuro}/m^{3}$ ]
Masterplan Files
The masterplan for each competition stage should be prepared using one of the template files provided in the supplementary materials. All three file formats are accepted to provide a trade-off between flexibility and usability.
The Excel file is the most user-friendly way to describe the solution. Each lever is described in a separate sheet and requires adding one entry per row and filling in the columns. The advantage is that most options can simply be selected using the dropdown menus in the columns. The downside is that it doesn’t offer the flexibility to produce custom policies for budget allocation (Section 2.3.1.1) and NRW interventions (Section 2.3.1.2).
The YAML format describes the masterplan using the same structure we introduce for all the levers outlined in Section 2.3. It allows users to define custom policies and to populate the file programmatically. However, preparing it may be somewhat more difficult as it needs to follow the specific structure shown in Listing 1. The masterplan can be provided in either JSON or YAML format as long as the structure remains unchanged.
Listing 1 Structure of the masterplan in YAML.
years: - year: YYYY national_policies: # Amend here budget allocation and other national policies national_interventions: # Install here inter-provincial pipes connecting water utilities water_utilities: - water_utility: WUxx policies: # Amend here policies for this water utility interventions: # Report here interventions for this water utility (e.g., pipe installations)